### v1

# Why machine learning models fail to fully capture epistemic uncertainty

Abstract
In recent years various supervised learning methods that disentangle aleatoric and epistemic uncertainty based on second-order distributions have been proposed. We argue that these methods fail to capture critical components of epistemic uncertainty, particularly due to the often-neglected component of model bias. To show this, we make use of a more fine-grained taxonomy of epistemic uncertainty sources in machine learning models, and analyse how the classical bias-variance decomposition of the expected prediction error can be decomposed into different parts reflecting these uncertainties. By using a simulation-based evaluation protocol which encompasses epistemic uncertainty due to both procedural- and data-driven uncertainty components, we illustrate that current methods rarely capture the full spectrum of epistemic uncertainty. Through theoretical insights and synthetic experiments, we show that high model bias can lead to misleadingly low estimates of epistemic uncertainty, and common second-order uncertainty quantification methods systematically blur bias-induced errors into aleatoric estimates, thereby underrepresenting epistemic uncertainty. Our findings underscore that meaningful aleatoric estimates are feasible only if all relevant sources of epistemic uncertainty are properly represented.
footnotetext: *Equal contribution.footnotetext: 1Department of Data Analysis and Mathematical Modeling.
1 Introduction
Uncertainty estimation in machine learning has emerged as an important tool to understand the reliability of a model’s predictions. In supervised learning, one often distinguishes between two types of uncertainty: aleatoric and epistemic uncertainty [1, 2]. Assuming that the ground truth data generating process can be described by a conditional probability distribution $p(y|\boldsymbol{x})$, aleatoric uncertainty reflects the inherent randomness of it, and is hence irreducible. Epistemic uncertainty, on the other hand, appears because the unknown $p(y|\boldsymbol{x})$ needs to be estimated using data. As a result, epistemic uncertainty can in principle be reduced by collecting more data, improving model specification, or changing the learning algorithm. While the correct disentanglement [3, 4] of these two types of uncertainty is crucial not only for model interpretability, but also for tasks like out-of-distribution detection [5] and active learning [6], it is far from trivial. Recent works show that using common methods of disentanglement in classification like the information-theoretic decomposition [3] not only is theoretically questionable [7], but can also lead to almost perfectly correlated aleatoric and epistemic uncertainty estimates [8]. In regression, common disentanglement methods like the variance-based decomposition [9, 3, 10] are often used in practice [9, 11, 12], but a thorough theoretical analysis of how well aleatoric and epistemic uncertainty are represented is still missing. Our main goal in this study is therefore to raise three main concerns:
Different sources of epistemic uncertainty make its evaluation troublesome. Epistemic uncertainty arises from various sources: Distribution shifts between training and test data [13], a limited training sample, an ill-defined model or procedural randomness [14] in the training algorithm. The ways to evaluate representations of uncertainty are equally vast, where the most common ones are arguably via the performance on downstream tasks [5, 6, 15]. A more theoretical line of work proposes axioms [7] for evaluating second-order distributions, stating several criteria that a second-order distribution has to fulfill in order to faithfully represent epistemic uncertainty. Recently, these axioms were adapted for the regression case [16], highlighting differences between entropy-based and variance-based disentanglement. A more pragmatic approach is the usage of a reference distribution [17] or a gold standard [18], allowing for the relative comparison of second-order distributions. Despite a variety of existing evaluation criteria, we argue that they often only evaluate one specific source of epistemic uncertainty. The aim of out-of-distribution detection methods is typically to capture uncertainty due to a shift between training and test data, while methods under the i.i.d. setting often implicitly assume that any lack of confidence in model parameters arises solely from a limited data sample, neglecting model bias or procedural randomness [14].
It is questionable which source of epistemic uncertainty existing methods capture. Various sources contribute to epistemic uncertainty [19], ranging from having limited training data, a wrongly defined model class and variability in the learning algorithm, to distribution shifts between training and test data. To account for the uncertainty in $p(y|\boldsymbol{x})$, epistemic uncertainty is often represented via second-order distributions over the parameters of the (first-order) distribution $p(y|\boldsymbol{x})$. This second-order distribution assigns probabilities to all candidate first-order probabilities $\hat{p}(y|\boldsymbol{x})$, thereby quantifying how much the aleatoric uncertainty estimate can be trusted. Common ways to learn such a second-order representation include Bayesian neural networks [20, 21], Gaussian processes [22, 23, 24], ensemble models [9] or evidential deep learning [12, 25, 13]. However, it is still unclear which sources of epistemic uncertainty the respective methods capture. Despite the common view that epistemic uncertainty reflects all “lack of knowledge,” practical methods often overlook important contributors, such as biases in the hypothesis space, and the difference between data- and procedural uncertainty [14]. The misinterpretation of which part of the epistemic uncertainty is being captured by common uncertainty quantification methods not only challenges the reliability of their usage for downstream tasks, but also the meaningfulness of uncertainty disentanglement.
Figure 1: Case where a high model bias is falsely attributed to aleatoric uncertainty: Due to limited data and complexity of the generating function, the underlying epistemic uncertainty to the left side of the $x$-axis is high. The model (falsely) predicts it as high aleatoric uncertainty. See Sec.4 for details.
Our main concern is that correct aleatoric uncertainty estimation is only possible if epistemic uncertainty is sufficiently low. While aleatoric uncertainty arises from the irreducible randomness in the true conditional distribution $p(y|\boldsymbol{x})$, its estimate $\hat{p}(y|\boldsymbol{x})$ depends on how well the model captures the underlying data generating process. Specifically, it is not reliably quantifiable if epistemic uncertainty remains high. If a model has little information about the underlying data generating function, its systematic errors can wrongly be attributed to measurement noise, thereby overestimating aleatoric uncertainty. Conversely, too low an estimate of epistemic uncertainty can lead to false trust in the assessed aleatoric uncertainty. Figure 1 shows an exemplary setting where a Gaussian process misclassified high epistemic uncertainty as high aleatoric uncertainty, due to high model bias.
Position. We call for a more fine-grained assessment of disentanglement of aleatoric and epistemic uncertainty. While it is crucial to distinguish between these two sources, one must identify which portion of the epistemic uncertainty a given model truly represents. Furthermore, we argue that one can only make meaningful statements about aleatoric uncertainty once epistemic uncertainty is correctly estimated as sufficiently low.
Contribution. To underline our concerns, we conduct an analysis of epistemic and aleatoric uncertainty estimation for second-order uncertainty representations. Focusing on the regression setting, we examine different sources of epistemic uncertainty, thereby including the often neglected procedural uncertainty, and derive theoretical expressions for them by usage of the classical bias-variance decomposition (Section 2). Subsequently, we discuss how epistemic uncertainty can be evaluated in a correct manner using statistical simulations, and we show with an illustrative example that the underlying model bias can lead to untrustworthy estimates if not taken into account, even if parts of the epistemic uncertainty are correctly estimated (Section 3). Furthermore, by assessing bias, data- and procedural uncertainty conceptually and experimentally, we illustrate that common second-order uncertainty quantification methods only capture parts of the epistemic uncertainty, often confusing it with aleatoric uncertainty (Section 4).
2 Towards a fine-grained definition of epistemic uncertainty
For the scope of this paper, we consider a heteroscedastic regression setting with instance space $\mathcal{X}\subset\mathbb{R}^{D}$, label space $\mathcal{Y}\subset\mathbb{R}$, and a dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ of size $N\in\mathbb{N}$, where each $(\boldsymbol{x}_{i},y_{i})$ is an independent and identically distributed (i.i.d.) realization of an unknown probability distribution $P$ on $\mathcal{X}\times\mathcal{Y}$. Let $p(y|\boldsymbol{x})$ be the associated conditional probability to observe label $y$ as an outcome given $\boldsymbol{x}$. In what follows we consider the following data-generating process:
$y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i},$ (1)
where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$. Assuming a conditional Gaussian distribution is a limiting assumption, but it has no substantial influence on the position we take in this paper. For a given $\boldsymbol{x}\in\mathcal{X}$, we define $\boldsymbol{\theta}(\boldsymbol{x})=(f(\boldsymbol{x}),\sigma^{2}(\boldsymbol{x}))$ as the parameters of the underlying normal distribution. One common representation of uncertainty is via a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ over these first-order parameters. Measures such as variance or entropy can then be used to estimate aleatoric and epistemic uncertainty. An estimate $\hat{p}(y|\boldsymbol{x})$ of the predictive distribution as well as the prediction $\hat{y}$ can then be obtained by Bayesian model averaging as
[…]
2.1 The difficulty of defining epistemic uncertainty
Figure 2: Sources of epistemic uncertainty in machine learning, where $p(y|\boldsymbol{x})$ denotes the true data-generating process. See text for further details. Figure adapted from Huang et al. [14].
Unlike aleatoric uncertainty, epistemic uncertainty does not have a universally accepted ground-truth value. Instead, it is inherently context dependent and must always be viewed relative to some benchmark. On the other hand, for different representations one can derive certain logical criteria that the aleatoric and epistemic uncertainty estimates have to fulfill. Empirically, Bengs et al. [26] argued that a second-order distribution should ideally be monotonous with respect to sample size and converge to a Dirac delta distribution if the dataset size goes to infinity. However, we argue that especially these empirical criteria are far from being fully established yet, and miss certain sources of epistemic uncertainty in practical applications, like procedural uncertainty [14], i.e. uncertainty due to randomness in the learning process. To this end, we provide a thorough taxonomy of the different sources of epistemic uncertainty in the next section.
As aleatoric uncertainty corresponds to the intrinsic randomness of the data generating process, one can directly associate it with the stochastic term $\epsilon$ in (1), whose distribution in turn depends on the variance $\sigma^{2}$. Epistemic uncertainty, on the other hand, measures "everything else" that is in principle reducible by collecting more knowledge, hence all uncertainty with respect to the parameters $\boldsymbol{\theta}$, but also the chosen hypothesis space $\mathcal{H}$. This uncertainty can stem from different sources, and several authors have discussed some of these sources [2, 14, 13]. Unifying those perspectives, we will distinguish in total four different sources of epistemic uncertainty (see Figure 2 for an overview). A first important subdivision into three components can be made as follows:
  • Model uncertainty: arises when the hypothesis space $\mathcal{H}$ does not contain the true underlying distribution, i.e. $p(y|\boldsymbol{x})\mathrel{\ooalign{\hfil$\in$\hfil\cr\hfil$/$\hfil\cr}}\mathcal{H}$. In Figure 2 the notation $\hat{p}(y|\boldsymbol{x})$ is used to denote the Bayes-optimal model for a given hypothesis space. It can be formally defined as follows:
    $\hat{p}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{y|\boldsymbol{x}}\ell(y,\hat{p}),$
    where $\ell(\cdot,\cdot)$ can be any loss function for regression, such as negative log-likelihood.
  • Estimation uncertainty: quantifies how well the learning algorithm can approximate $\hat{p}(y|\boldsymbol{x})$, which is the best possible model given the hypothesis space.
  • Distributional uncertainty: is due to a shift in distribution between training and test data [13], respectively, and their associated conditional distributions $p(y|\boldsymbol{x})$ and $\tilde{p}(y|\boldsymbol{x})$.
Following recent work of Huang et al. [14], we further divide estimation uncertainty into data uncertainty and procedural uncertainty. To this end, we define $\mathcal{D}_{N}\sim P^{N}$ as the random variable reflecting the data uncertainty, where $P^{N}$ is the distribution over all datasets of size $N$. Similarly, we define $\gamma\sim P^{\gamma}$ as the random variable reflecting all procedural parameters (e.g. random weight initialization in Deep Ensembles), where $P^{\gamma}$ denotes the distribution over which these parameters are sampled. For a fixed dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$, data uncertainty can then be formally defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and
$\hat{p}_{N}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\gamma}\big{[}\frac{1}{N}\sum_{i=1}^{N}\ell(y_{i},\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}))\big{]}\,,$
where $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ denotes a model trained on a given dataset of size $N$ and procedural parameters $\gamma$. Similarly, for a fixed set of procedural parameters, procedural uncertainty can be formally defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and
$\hat{p}_{\gamma}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\mathcal{D}_{N}}\big{[}\frac{1}{N}\sum_{i=1}^{N}\ell(y_{i},\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}))\big{]}\,,$
Using this notation, estimation uncertainty is then defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$, as visualized in Figure 2.
2.2 Uncertainty from a bias-variance perspective
From a statisticians viewpoint, the classical bias-variance decomposition provides further insight into the interplay between estimation uncertainty and aleatoric uncertainty [24, 19]. Here we extend it to incorporate not only the data uncertainty induced by $\mathcal{D}_{N}$, but also the uncertainty from the procedural uncertainty given by $\gamma$.
$\mathbb{E}_{y|\boldsymbol{x}}\big{[}\mathbb{E}_{\mathcal{D}_{N},\gamma}(y-\hat{y})^{2}|\boldsymbol{x}\big{]}$ $=\mathbb{E}_{y|\boldsymbol{x}}\big{[}\mathbb{E}_{\mathcal{D}_{N},\gamma}(y-\mathbb{E}_{y|\boldsymbol{x}}[y|\boldsymbol{x}])^{2}|\boldsymbol{x}\big{]}+\mathbb{E}_{y|\boldsymbol{x}}\big{[}\mathbb{E}_{\mathcal{D}_{N},\gamma}(\mathbb{E}_{y|\boldsymbol{x}}\big{[}y|\boldsymbol{x}]-\hat{y})^{2}|\boldsymbol{x}\big{]}$
$=\sigma^{2}(\boldsymbol{x})+\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})+\text{bias}^{2}\big{[}\hat{y}|\boldsymbol{x}\big{]}.$ (3)
The first term in Eqn. 2.2 corresponds to the intrinsic variability within the data-generating process, and hence to aleatoric uncertainty. The second term reflects the variance of the predictions $\hat{y}$ under data and procedural uncertainty, and can be further decomposed by means of the law of total variance as
$\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})=\underbrace{\mathbb{E}_{\mathcal{D}_{N}}\big{[}\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big{]}}_{\text{procedural uncertainty}}+\underbrace{\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N})}_{\text{data uncertainty}}.$ (4)
The third term reflects the squared bias, which is the difference between the expected prediction $\hat{y}$ under data and model uncertainty and the true conditional expectation $\mathbb{E}_{y|\boldsymbol{x}}\big{[}y|\boldsymbol{x}\big{]}$. For universal approximators, such as many types of deep neural networks, the model uncertainty is zero [27, 28, 29]. Such models are asymptotically unbiased, but in the finite data regime, a bias is unavoidable to make learning possible [24]. For more restrictive hypothesis spaces, the bias even does not disappear when the training dataset becomes infinitely large. Thus, in both cases, the bias term cannot be ignored when analysing epistemic uncertainty.
3 Evaluating epistemic uncertainty via statistical simulations
Quantifying epistemic uncertainty objectively is challenging, and evaluating the correctness of such quantifications is equally difficult. While calibration tests exist for aleatoric uncertainty in probabilistic models [30], epistemic uncertainty lacks established evaluation protocols and is often assessed indirectly through downstream tasks like out-of-distribution detection [5, 15, 31], robustness to adversarial attacks [32, 33], and active learning [34]. We believe that correctness of second-order epistemic uncertainty representations can only be evaluated through statistical simulations, in which bias and variance can be analysed individually. To this end, Jürgens et al. [17] recently introduced the notion of a reference distribution, which is the theoretical distribution over first-order models if training data could be resampled. Here we further extend this notion by taking the variance due to procedural uncertainty into account:
Definition 3.1.
Consider a hypothesis space of models with $\phi\in\Phi$ model parameters and $\gamma\in\Gamma$ procedural parameters. We define the reference distribution w.r.t. $P^{N}$ and $P^{\gamma}$ as
[…]
where $\boldsymbol{\hat{\theta}}_{N,\gamma}$ is the empirical estimate of $\boldsymbol{\theta}$, obtained via empirical risk minimization.
Intuitively, the reference distribution describes the uncertainty in the first-order parameters, induced by having a limited dataset and procedural uncertainty due to different weight initializations. In simulation studies, one can estimate the reference distribution by drawing $n_{d}$ training datasets $\mathcal{D}_{N_{1}},\dots,\mathcal{D}_{N_{n_{d}}}$ and $n_{\gamma}$ procedural parameters $\gamma_{1},\dots,\gamma_{n_{\gamma}}$ from $P^{N}$ and $P^{\gamma}$, respectively, and calculating the empirical distribution function.
The variance of the reference distribution describes the variance due to limited data expressivity, as well as the procedural variance, and can be decomposed by means of the law of total variance in Eqn. 4. In addition, in statistical simulations, where the data-generating process $p(y|\boldsymbol{x})$ is known, the bias of the reference distribution can also be measured.
Example 3.2 (An illustrative statistical simulation).
We illustrate this argument by adapting the synthetic regression setting from Rossellini et al. [35]. Precisely, we generate datasets $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\sim Beta(1.2,0.5)$, $y_{i}=\sin(\frac{1}{5(x_{i}+0.16)^{3}})+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(x_{i}))$ with $\sigma^{2}(x)=x^{4}$. The setup can be seen in Figure 1. This dataset highlights the differences between aleatoric and epistemic uncertainty: for low $x$ values, the complexity of the underlying data-generating function is high, while data is limited, leading to high epistemic uncertainty. On the other hand, for $x$ close to $1$, aleatoric uncertainty is high, while simultaneously having more data points and a less complex underlying function, implying low epistemic uncertainty. In order to estimate the reference distribution we train models with $n_{d}$ training datasets and $n_{\gamma}$ random weight initializations. As the underlying model class for obtaining the individual risk minimizers $\boldsymbol{\hat{\theta}}_{\gamma,N}$, we use an MLP. Details about the experimental setup are discussed in Appendix A.1111The implementation for all our examples can be found in this GitHub repository.. Individual estimates for data- and procedural uncertainty can be obtained by the decomposition formula in Eqn.4. The bias can be calculated as the absolute distance between the mean prediction and the underlying ground truth conditional expectation.
Figure 3: Estimated procedural and data uncertainty of the reference distribution for increasing sample sizes. The reference distribution is estimated by resampling training data and procedural parameters, using $n_{\gamma}=10$ and $n_{d}=20$. Procedural and data uncertainty are calculated using the decomposition formula in Eqn. 4. Even when the model predictions (complete line) have an important deviation (bias) from the ground-truth expected value (dashed line), the epistemic uncertainty estimate decreases with more data.
Figure 3 shows the obtained reference distribution for different training dataset sizes. First, we see that the total variance of the reference distribution decreases when sample size increases, supporting the claim that the reference distribution properly represents data uncertainty. Second, we see that the estimate of the procedural uncertainty also decreases with an increasing sample size, which is not unexpected, because initial procedural configurations become less important when more data is observed. Third, in the region with high epistemic uncertainty, the reference distribution has a very large bias, due to the underlying complexity of the data generating function. The bias is in fact so high that aleatoric uncertainty cannot be correctly estimated anymore This last finding confirms our central position: Any claim to properly represent epistemic uncertainty must examine not only how different sources of uncertainty are modeled, but also whether bias is systematically distorting the partitioning of uncertainty. When the bias is high, estimates of both epistemic and aleatoric uncertainty can be completely misleading.
4 Which part of epistemic uncertainty do current methods capture?
Departing from a Bayesian perspective, we first describe in this section how various second-order methods claim to obtain disentangled estimates for aleatoric and epistemic uncertainty. Secondly, we analyse two representative methods that approximate the posterior distribution over the parameters in more detail, illustrating that they tackle different sources of epistemic uncertainty. In the last part of this section, we present experimental results for eight second-order methods, highlighting that all of them have difficulties to model the different sources of epistemic uncertainty in a complete manner.
4.1 Second-order uncertainty quantification methods
Bayesian methods form principled approach to quantify and interpret epistemic uncertainty by modelling a prior and posterior distribution over the model parameters and computing the second-order distribution via Eqn. 2. For the regression setup that we consider in this paper, a common approach is to use Heteroscedastic Gaussian Processes (GP, [23, 36]), which place priors on the parameters $f(\boldsymbol{x})$ and $\sigma^{2}(\boldsymbol{x})$, thereby analytically computing the predictive distribution using two Gaussian processes, $f(\boldsymbol{x}),\log\sigma^{2}(\boldsymbol{x})\sim GPs$. For Bayesian neural networks, which have gained popularity in recent years, the posterior and second-order distribution cannot be analytically computed, due to the intractability of $p(\boldsymbol{\phi}|\mathcal{D}_{N})\propto p(\mathcal{D}_{N}|\boldsymbol{\phi})p(\boldsymbol{\phi})$ induced by the hypothesis space [1]. Due to the impossibility of performing exact inference in Bayesian neural networks, one typically obtains $d$ different samples $\{\boldsymbol{\phi}_{j}\}_{j=1}^{d}$ from $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ by use of approximation techniques that aim to find a posterior distribution via sampling methods. Here, the second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ of Eqn. 2 is
is then approximated by Monte Carlo integration:
$q(\boldsymbol{\theta}|\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\phi}\sim p(\boldsymbol{\phi}|\boldsymbol{x})}\big{[}q(\boldsymbol{\theta}\mid\boldsymbol{x},\boldsymbol{\phi})\big{]}\approx\frac{1}{d}\sum_{j=1}^{d}q(\boldsymbol{\theta}\mid\boldsymbol{x},\boldsymbol{\phi}_{j}).$ (6)
In the experiments below, we employ three of the most common methods to obtain $\{\boldsymbol{\phi}_{i}\}_{j=1}^{M}$. Variational inference, (VI, [37, 38]) considers a variational (simpler) distribution $q(\boldsymbol{\phi})$ and optimizes the lower bound of the marginal likelihood (ELBO). Hamiltonian Monte Carlo (HMC) uses Hamiltonian dynamics to sample $\{\boldsymbol{\phi}_{j}\}$ from the true posterior. Laplace approximation (LA, [39]) builds a Taylor expansion around the maximum-a-posteriori estimate $\boldsymbol{\phi}_{\text{MAP}}$, yielding a Gaussian approximation.
Ensemble methods are also often interpreted as performing approximate Bayesian inference by finding multiple modes of $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ after training multiple independent models. [40, 41]. In this work, we consider three types of ensembles. Deep Ensembles (DE, [9]), train $d$ networks on the same dataset $\mathcal{D}_{N}$, but with different random initialization of the weights $\boldsymbol{\phi}$. Similarly, in Bootstrap Ensembles (BE), $M$ networks are trained using different weight initializations, but each network uses a bootstrap sample $\mathcal{D}^{(j)}\subseteq\mathcal{D}_{N}$. While Bootstrap Ensembles have been shown to perform worse than training $d$ networks using Deep Ensembles [9], we include them in our analysis to estimate the effect of data uncertainty on the epistemic uncertainty estimates. Monte Carlo Dropout (MCD, [42]) can also be understood as a Bayesian approximation, where one trains a single neural network with dropout rate $\rho$ and performs multiple stochastic forward passes at inference time with the same dropout rate. In a different direction, evidential deep learning methods learn directly the second-order distribution over the predictive parameters $\theta$. Deep evidential regression (DEL, [12]) learns hyperparameters of a conjugate prior (e.g. Normal-Inverse-Gamma for the univariate normal distribution) directly using a single network trained by second-order loss minimization.
From a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$, one can decompose its representation of uncertainty into an estimate of aleatoric and epistemic uncertainty. Proposed measures focus on entropy [3], distance to reference sets [10], or proper scoring rules [43]. However, the most commonly used one for regression is arguably the variance-based decomposition [3, 10], which considers the variance of the prediction $\hat{y}$, obtained via model averaging as in Eqn. 2. The latter can be decomposed into (estimates of) an aleatoric and an epistemic part:
$\text{Var}(\hat{y}|\boldsymbol{x})=\underbrace{\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta}\mid\boldsymbol{x})}\big{[}\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big{]}}_{\text{ aleatoric estimate}}+\underbrace{\text{Var}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta}\mid\boldsymbol{x})}\big{(}\mathbb{E}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big{)}}_{\text{epistemic estimate}}$ (7)
Here, the first term estimating aleatoric uncertainty describes the average predicted variance $\sigma^{2}$. The second term captures the variance of the mean prediction $f$ and hence an estimate of epistemic uncertainty. Note here that, unlike in the bias-variance decomposition in Eqn. 2.2, the bias in the prediction is not represented by this decomposition. Since any second-order distribution does not incorporate ground-truth knowledge, any systematic bias goes unaccounted. Moreover, second-order uncertainty quantification methods lack informativeness about which source of epistemic uncertainty they capture. In the next paragraph we illustrate these two concerns for two methods that serve as representative examples.
4.2 Two representative cases
Example 4.1 (Deep Ensembles purely capture procedural uncertainty).
Let $\gamma$ denote the random seed (weight initialization, data ordering etc.) that differentiates each ensemble member’s training, which in turn are trained on the same dataset $\mathcal{D}_{N}$. For a Deep Ensemble of size $d$, each ensemble member’s predictive parameters are indexed over the set of random seeds $\{\gamma_{1},\dots,\gamma_{d}\}$, obtaining a set of predictions $\{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})=(f_{\gamma_{i}}(\boldsymbol{x}),\sigma_{\gamma_{i}}^{2}(\boldsymbol{x}))\}_{i=1}^{d}$. The empirical second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ can be approximated as a mixture of equally weighted Dirac measures
$q(\boldsymbol{\theta}|\boldsymbol{x})=\frac{1}{d}\sum_{i=1}^{d}\delta_{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})}(\boldsymbol{\theta}),$
[…]
$\text{Var}(\hat{y}|\boldsymbol{x})$ $=\frac{1}{d}\sum_{i=1}^{d}\sigma_{\gamma_{i}}^{2}(\boldsymbol{x})+\frac{1}{d}\sum_{i=1}^{d}(f_{\gamma_{i}}(\boldsymbol{x})-\hat{f}(\boldsymbol{x}))^{2}$
$\longrightarrow$ $\,\mathbb{E}_{\gamma}\big{[}\text{Var}(\hat{y}|\boldsymbol{x},\gamma)\big{]}+\text{Var}_{\gamma}\big{(}\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big{)},\,d\rightarrow\infty.$
Here, the first part corresponds to the aleatoric uncertainty estimate and the epistemic uncertainty estimate is given by
$\text{Var}_{\gamma}\big{(}\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big{)}=\mathbb{E}_{\gamma}\big{[}(\mathbb{E}_{x|y}[\hat{y}|\boldsymbol{x},\gamma]-\mathbb{E}_{\gamma}[\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]])^{2}\big{]}=\mathbb{E}_{\gamma}\big{[}(\hat{y}_{N,\gamma}(\boldsymbol{x})-\hat{y}_{N}(\boldsymbol{x}))^{2}\big{]},$
where $\hat{y}_{N,\gamma}(\boldsymbol{x}):=f_{\gamma}(\boldsymbol{x})$ corresponds to the prediction of an individual ensemble member trained with random initialization $\gamma$ on $\mathcal{D}_{N}$ and $\hat{y}_{N}(\boldsymbol{x}):=\hat{f}(\boldsymbol{x})$ of the aggregated ensembles. Hence, we see that the predicted epistemic uncertainty corresponds to the procedural uncertainty as illustrated in Figure 1 a). Model uncertainty through bias as well as data uncertainty are not represented.
While Deep Ensembles might explore multiple (local) optima of the training loss by using different random seeds [40], Bayesian approaches attempt to explicitly approximate the posterior over parameters, which can yield more principled estimates of epistemic uncertainty. However, the practical methods used to approximate the posterior distribution also face restrictions: Approximate inference methods like the Laplace approximation or variational inference often capture only a single local mode [40], employ simplifying assumptions [18], or can be computationally costly in high-dimensional weight spaces. Other issues arise with methods that try to estimate the second-order distribution directly via loss minimisation, such as Deep Evidential Regression.
Example 4.2 (DER may fail to provide to faithful epistemic uncertainty estimates).
Deep Evidential Regression [12] learns to predict the parameters of a Normal–Inverse-Gamma second–order distribution
$q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})=\mathrm{NIG}\!\bigl{(}\gamma(\boldsymbol{x}),\nu(\boldsymbol{x}),\alpha(\boldsymbol{x}),\beta(\boldsymbol{x})\bigr{)},$
which forms the conjugate prior of the normal distribution. The parameters $(\gamma,\nu,\beta,\alpha)$ are learned by minimising a second-order loss, i.e. the loss function one derives when doing Bayesian inference with the NIG prior, together with a regularization term called the evidence, which controls the uncertainty for data points far from the training set. Using the variance-based decomposition, one obtains analytical expressions for the aleatoric and epistemic uncertainty estimates, namely
$\mathbb{E}_{\boldsymbol{\theta}\sim q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})}[\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big{]}=\frac{\beta}{\alpha-1},\quad\text{Var}_{\boldsymbol{\theta}\sim q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})}\big{(}\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big{)}=\frac{\beta(\alpha-1)}{\nu}.$
However, a line of work has shown that the method has severe drawbacks in the optimization [26, 44], as well as the estimation of aleatoric and epistemic uncertainty [45], and the representation of the second-order distribution [17], making it more of a heuristic than a principled method.
4.3 Synthetic data illustration
To illustrate the limitations of existing uncertainty quantification methods in producing reliable disentangled estimates, we showcase the behavior of the eight methods described in Section 4.1 for the problem setting described in Example 3.2. This section focuses primarily on the results obtained using Deep Ensembles, but we emphasize that the results for the other seven methods are also consistent with Deep Ensembles, and coherent with our argumentation. Results for the remaining seven methods are provided in Appendix B, and detailed implementation specifications are included in Appendix A.2.
Figure 4: Deep Ensemble estimates, obtained with different sample sizes $N\in\{50,100,500\}$. The top row shows the training data points (blue crosses), the true conditional mean (dashed line), the model predictions (complete line), the $95\%$ confidence interval for aleatoric uncertainty (purple), as well as the true $95\%$ confidence interval of the aleatoric uncertainty (gray). The bottom row shows the estimate of epistemic uncertainty (green).
Figure 4 shows a detailed illustration of the results for Deep Ensembles. Although the aleatoric uncertainty estimates on the right-hand side of the data support improve as the sample size increases, the behaviour on the left-hand side remains problematic. Specifically, the model persistently overestimates aleatoric uncertainty. A plausible explanation is that the approximate symmetry of the observations around the learned conditional mean induces the model to misinterpret systematic deviation as inherent data noise. Consequently, this misestimation is not corrected with additional data; instead, the model becomes increasingly confident in its erroneous predictions and continues to interpret the discrepancy from the ground-truth conditional mean as aleatoric variability. By contrast, epistemic uncertainty decreases monotonically with sample size, mirroring the behaviour observed for the reference distribution. The often-repeated claim in the literature that aleatoric uncertainty estimates can be trusted once epistemic uncertainty is low is therefore not correct: when the training set is sufficiently large, in a difficult-to-learn setting, the model remains overconfident in highly biased predictions, producing inflated aleatoric uncertainty estimates. These observations support our two principal conclusions. First, current measures of epistemic uncertainty and existing disentanglement frameworks fail to capture systematic bias, thereby underestimating epistemic uncertainty. Second, the resulting aleatoric-uncertainty estimates misclassify epistemic uncertainty with data noise and hence do not achieve a precise disentanglement of the two sources of uncertainty. Bias is approximated as specified in Eqn. 2.2: datasets are repeatedly resampled from the true data-generating process, models are trained on each sample, and the mean deviation of the predictions from the true conditional mean is computed. Comprehensive results for every method are presented in Table 1; the table shows that bias is frequently larger than—and largely uncorrelated with—the corresponding epistemic uncertainty estimates. These findings underscore our central position: Any attempt to disentangle aleatoric from epistemic uncertainty must assess not only how each component is modelled, but also the extent to which systematic bias distorts their partitioning.
The results for the remaining methods are provided in Appendix A. In this section, we highlight two additional key observations. First, we observe that Bayesian approaches—specifically Laplace Approximation (LA) and Hamiltonian Monte Carlo (HMC)—require a larger amount of data to produce accurate estimates of aleatoric uncertainty in the right-hand region of the input space. This behavior aligns with theoretical expectations of Bayesian inference: the increased data requirement reflects the greater difficulty these models encounter in learning the underlying data-generating function, which in turn leads to higher epistemic uncertainty—consistent with models that make inaccurate predictions. Second, the Deep Evidential Regression (DER) method consistently produces poor estimates of aleatoric uncertainty across all sample sizes and throughout the entire input domain. These inaccuracies are accompanied by overconfident predictions, as evidenced by the systematic underestimation of epistemic uncertainty. This flawed characterization of uncertainty is consistent with findings reported in recent literature [17].
5 Discussion
In this work, we address the problem of blurriness in the definition of epistemic uncertainty for machine learning models and the difficulties that come with it. In attempting to solve this problem, we first unified different definitions of epistemic uncertainties for machine learning models into a formal taxonomy. In our illustrative simulations, we work within a heteroscedastic regression framework; nevertheless, the key findings are model-independent and hence translate to classification problems. Using a statistical sampling-based evaluation protocol, we showed that even if parts of the arising epistemic uncertainty are well captured, the underlying model bias can still lead to an overall underestimate of epistemic uncertainty. By differentiating between procedural and data uncertainty, we find that different uncertainty estimation methods merely capture parts of the remaining epistemic uncertainty. Furthermore, the non-assessment of bias seemingly leads to an artificially high estimate of aleatoric uncertainty. Our call is therefore not only for deriving more task-specific and fine-grained evaluation protocols of epistemic uncertainty in practical applications, but also a proper analysis of model bias when estimating it. It might be argued that the goal of an epistemic uncertainty estimate is not to quantify bias. However, precisely by not doing so, the disentanglement of uncertainties becomes problematic.
Appendix A Experimental details
A.1 Reference distribution approximation
To obtain an empirical approximation of the reference distribution, we sample $20$ different datasets. For each dataset, we train $10$ different MLP models with different random weight initialization. Each MLP uses the whole dataset. The MLP models have $4$ hidden layers with $100$ neurons each. We use a learning rate of 0.009 and train the models for $500$ epochs. We repeated the process for different training dataset sizes: $50,100,500$ with batch sizes $32,32,64$ respectively. Seeds for the datasets are $\{7,42,2024,123,999,50,100,150,200,250,300,350,400,450,500,550,600,650,700,750\}$.
A.2 Experimental details by method
All methods are evaluated for different training dataset sizes: $\{50,100,500\}$ with batch sizes $\{32,32,64\}$ respectively. The experiments are repeated for $4$ different seeds: $7,42,123,999$ and $2024$. To conduct the experiments, a hyperparameter tuning over relevant parameters was performed.
Deep Ensemble. Trained with an MLP of $4$ hidden layers with $100$ neurons each, learning rate of 0.009. Training epochs: $50,250,500$. Every member of the deep ensemble uses the whole dataset.
Bootstrap Ensemble. Trained as the Deep Ensemble, but each ensemble member uses a boostrap sample with $60\%$ of the training data.
Monte Carlo Dropout. A single MLP of $4$ hidden layers with $100$ neurons each is trained. A dropout rate of $0.2$ is used at training and inference time. A learning rate of 0.002 is used.
Variational Inference. This model is trained based on Blundell et al. [46]. An MLP of $4$ hidden layers with $100$ neurons each converted to a Bayesian neural network is used. The model is trained with a learning rate of $0.005$, $200$ burn-in epochs, a beta value of $5$, $10$ Monte Carlo samples at training time, and $500$ at test. The model is evaluated for $250$ and $500$ training epochs.
Laplace Approximation. An MLP of $4$ hidden layers with $100$ neurons each is used to find the MAP estimate, together with a learning rate of $0.005$. For the Laplace step, a prior precision of $1$ and a noise value of $2$ are used. The model is trained for $200,400$, and $800$ epochs, and uses $1000$ samples from the posterior distribution. When using a training sample size of $500$, a batch size of 128 is used.
Hamiltonian Monte Carlo. An MLP of $4$ hidden layers with $100$ neurons each is used as the underlying model. The sampling step consists of $200$ samples with a sampling step of $0.00015$. The first $50$ samples are always burnt. A $\tau$ value of $1$ is used as the indicator for a bendy function. At inference time, we sample $1000$ times from the posterior distribution and drop the first $50$.The model is trained for $1000,2000$, and $3000$ epochs. When using a sample size of 500, a batch size of 128 is used.
Deep Evidential Regression. Leaning on the setup of Amini et al. [12], an MLP of $4$ hidden layers with $100$ neurons each is used as the underlying model. The model is trained with a learning rate of $0.0003$ over $5000$ epochs, and a regularization strength of $\lambda=0.01$.
Heteroscedastic Gaussian Process. Leaning on the variational-sparse recipe of Titsias [36], we use the GPytorch ApproximateGP module to train a two-output GP, trained with an RBF kernel. The number of inducing points (see [24] for an explanation) is set to $256$, the learning rate to 0.005, epochs to $2000$.
Appendix B Further experimental results
In this section, we include further experimental results for the other methods that we tested. The data is generated according to Example 3.2. We conduct the same experiment as described for the Deep Ensemble case in Section 4.3. The hyperparameters of each method were chosen according to Section A. The results can be seen in Figure 5-11. We see that every method with a low epistemic uncertainty estimate overestimates the aleatoric uncertainty in the left-hand side of the figure due to their inability to acknowledge the bias as part of their epistemic uncertainty representations. Also, for more data-intensive methods, such as the Bayesian approaches, a more careful and realistic estimate of the epistemic uncertainty is obtained.
Figure 5: VI estimates, obtained with different sample sizes as in Figure 4.
Figure 6: GP estimates, obtained with different sample sizes as in Figure 4.
Figure 7: Bootstraped ensemble estimates, obtained with different sample sizes as in Figure 4.
Figure 8: MC Dropout estimates, obtained with different sample sizes as in Figure 4.
Figure 9: HMC estimates, obtained with different sample sizes as in Figure 4.
Figure 10: Laplace estimates, obtained with different sample sizes as in Figure 4.
Figure 11: DER estimates, obtained with different sample sizes as in Figure 4.
Table 1: Aleatoric and epistemic uncertainty estimates, bias, and distance from the aleatoric uncertainty to ground truth noise across models. Values are means over the test set from five runs; standard deviations are in parentheses.
Training sample size: N = 100 Model Aleatoric unc. Epistemic unc. Bias $\hat{\sigma}(x)-\sigma(x)$ Reference 0.83 (0.01) 0.041 (0.005) 0.084 (0.002) 0.248 (0.004) Bootstrap 0.87 (0.167) 0.006 (0.008) 0.125 (0.022) 0.239 (0.027) DeepEnsemble 0.87 (0.148) 0.007 (0.011) 0.147 (0.018) 0.235 (0.021) EDL 0.21 (0.002) 0.003 (0.0) 0.269 (0.01) 0.751 (0.005) ViELBONet 0.95 (0.064) 0.079 (0.001) 0.127 (0.021) 0.234 (0.022) MCDropoutNet 0.9 (0.097) 0.009 (0.002) 0.19 (0.026) 0.334 (0.033) HMCNet 9.08 (13.734) 2.131 (0.224) 0.179 (0.015) 7.87 (0.853) LaplaceNet 2.43 (5.559) 4.211 (3.047) 0.133 (0.024) 0.8 (0.54) GP 0.86 (0.048) 0.008 (0.001) 0.194 (0.013) 0.308 (0.02)
Training sample size: N = 500 Model Aleatoric unc. Epistemic unc. Bias $\hat{\sigma}(x)-\sigma(x)$ Reference 1.04 (0.016) 0.013 (0.001) 0.044 (0.001) 0.076 (0.003) Bootstrap 0.98 (0.1) 0.003 (0.004) 0.085 (0.001) 0.131 (0.01) DeepEnsemble 0.99 (0.136) 0.004 (0.005) 0.1 (0.002) 0.122 (0.01) EDL 0.67 (0.247) 0.013 (0.007) 0.168 (0.016) 0.967 (0.031) ViELBONet 1.06 (0.115) 0.009 (0.0) 0.097 (0.005) 0.122 (0.012) MCDropoutNet 1.06 (0.079) 0.008 (0.003) 0.119 (0.006) 0.163 (0.008) HMCNet 4.17 (20.356) 0.378 (0.017) 0.129 (0.003) 1.936 (1.495) LaplaceNet 1.29 (3.777) 0.487 (3.802) 0.115 (0.016) 0.187 (0.345) GP 1.04 (0.069) 0.008 (0.003) 0.172 (0.005) 0.152 (0.014)
In Table 1, we calculate the average aleatoric and epistemic uncertainty estimates for the eight different methods we analyse as showcases for our arguments. The last two columns show the average bias with respect to the ground truth conditional mean and the mean absolute distance from the aleatoric uncertainty estimate to the ground truth noise, respectively. It shows how methods that become overconfident with more training data are highly biased towards the true conditional mean. Furthermore, no correlation between epistemic uncertainty and bias is observed, showing how the unfaithfulness of the estimation is not only due to the order of magnitude of the estimands but also to the lack of information about the systematic deviations from the ground truth.