### v2

Abstract
In recent years various supervised learning methods that disentangle aleatoric and epistemic uncertainty based on second-order distributions have been proposed. We argue that these methods fail to capture critical components of epistemic uncertainty, particularly due to the mostly neglected component of model bias, but also due to only capturing procedural but not data uncertainty. By using a simulation-based evaluation protocol that encompasses epistemic uncertainty due to both procedural and data uncertainty components, we illustrate that current methods rarely capture the full spectrum of epistemic uncertainty. Through theoretical insights and experimental results on synthetic and real-world data, we find that high model bias can lead to misleadingly low estimates of epistemic uncertainty, and common second-order uncertainty quantification methods for regression seem to blur bias-induced errors into aleatoric estimates, thereby underrepresenting epistemic uncertainty. Our findings underscore that meaningful aleatoric estimates are feasible only if all relevant sources of epistemic uncertainty are properly represented.
keywords:
uncertainty estimation, uncertainty disentanglement, heteroscedastic regression
1 Introduction
Uncertainty estimation in machine learning has emerged as an important tool to understand the reliability of a model’s predictions. In supervised learning, one often distinguishes between two types of uncertainty: aleatoric and epistemic uncertainty [1, 2]. Assuming that the ground truth data generating process can be described by a conditional probability distribution $p(y|\boldsymbol{x})$, aleatoric uncertainty reflects the inherent randomness of it, and is hence irreducible. Epistemic uncertainty, on the other hand, appears because the unknown $p(y|\boldsymbol{x})$ needs to be estimated using data. As a result, epistemic uncertainty can in principle be reduced by collecting more data, improving model specification, or changing the learning algorithm. While the correct disentanglement [3, 4] of these two types of uncertainty is crucial not only for model interpretability, but also for tasks like out-of-distribution detection [5] and active learning [6], it is far from trivial. Recent works show that using common methods of disentanglement in classification like the information-theoretic decomposition [3] not only is theoretically questionable [7], but can also lead to almost perfectly correlated aleatoric and epistemic uncertainty estimates [8]. In regression, common disentanglement methods like the variance-based decomposition [9, 3, 10] are often used in practice [9, 11, 12], but a thorough theoretical analysis of how well aleatoric and epistemic uncertainty are represented is still missing. Our main goal in this study is therefore to raise three main concerns:
It is questionable which source of epistemic uncertainty existing methods capture. Epistemic uncertainty can arise from multiple sources [13], including distribution shifts between training and test data [14], limited training samples, ill-defined model classes, or procedural randomness in the training algorithm [15]. To represent these uncertainties, most approaches rely on second-order distributions over the parameters of the predictive (first-order) distribution $p(y|\boldsymbol{x})$. This second-order distribution assigns probabilities to all candidate first-order probabilities $\hat{p}(y|\boldsymbol{x})$, thereby quantifying how much the uncertainty estimate can be trusted. Common ways to learn such a second-order representation, include Bayesian neural networks [16, 17], Gaussian processes [18, 19, 20], ensemble models [9] or evidential deep learning [12, 21, 14]. However, it remains unclear which sources of epistemic uncertainty these methods actually capture. Most implicitly focus on the variability of model parameters, while overlooking systematic components such as model bias or hypothesis-space misspecification.
Different sources of epistemic uncertainty make its evaluation troublesome. The ways to evaluate representations of uncertainty are equally vast, where the most common ones are arguably via the performance on downstream tasks [5, 6, 22]. A more theoretical line of work proposes axioms [7] for evaluating second-order distributions, stating several criteria that a second-order distribution has to fulfill in order to faithfully represent epistemic uncertainty. Recently, these axioms were adapted for the regression case [23], highlighting differences between entropy-based and variance-based disentanglement. A more pragmatic approach is the usage of a reference distribution [24] or a gold standard [25], allowing for the relative comparison of second-order distributions. Despite a variety of existing evaluation criteria, we argue that they often only evaluate one specific source of epistemic uncertainty. The aim of out-of-distribution detection methods is typically to capture uncertainty due to a shift between training and test data, while methods under the i.i.d. setting often implicitly assume that any lack of confidence in model parameters arises solely from a limited data sample, neglecting model bias [26] or procedural randomness [15]. Even though the common view that epistemic uncertainty reflects all “lack of knowledge”, practical methods often overlook important contributors, such as biases in the hypothesis space, and the difference between data- and procedural uncertainty [15]. The misinterpretation of which part of the epistemic uncertainty is being captured by common uncertainty quantification methods not only challenges the reliability of their usage for downstream tasks, but also the meaningfulness of the disentanglement between epistemic and aleatoric.
6cm].49
6cm].49
Figure 1: (a): Synthetic experiment where a high model bias is falsely attributed to aleatoric uncertainty: Due to limited data points on the left side of the $x$-axis, and high complexity of the generating function, the underlying epistemic uncertainty there is high. The heteroscedastic Gaussian process model (falsely) predicts it as high aleatoric uncertainty. See Sec. 4 for details. (b): Sources of epistemic uncertainty in machine learning, where $p(y|\boldsymbol{x})$ denotes the true data-generating process. Different sources of epistemic uncertainty can lead to the estimate $\hat{p}(y|\boldsymbol{x})$ being further from the ground truth. See Sec. 2 for further details. Figure adapted from Huang et al. [15].
Our main concern is that correct aleatoric uncertainty estimation is only possible if epistemic uncertainty is sufficiently low. While aleatoric uncertainty arises from the irreducible randomness in the true conditional distribution $p(y|\boldsymbol{x})$, its estimate $\hat{p}(y|\boldsymbol{x})$ depends on how well the model captures the underlying data generating process. Specifically, it is not reliably quantifiable if epistemic uncertainty remains high. If a model has little information about the underlying data generating function, its systematic errors can wrongly be attributed to measurement noise, thereby overestimating aleatoric uncertainty. Conversely, a too low estimate of epistemic uncertainty can lead to false trust in the assessed aleatoric uncertainty. Figure 1(a) shows an exemplary setting where a Gaussian process misclassified high epistemic uncertainty as high aleatoric uncertainty, due to high model bias. Given these concerns, we call for a more fine-grained assessment of disentanglement of aleatoric and epistemic uncertainty. While it is crucial to distinguish between these two sources, one must identify which portion of the epistemic uncertainty a given model truly represents. Furthermore, we argue that one can only make meaningful statements about aleatoric uncertainty once epistemic uncertainty is correctly estimated as sufficiently low.
Contribution. To underline our concerns, we conduct an analysis of epistemic and aleatoric uncertainty estimation for second-order uncertainty representations. Focusing on the regression setting, we examine different sources of epistemic uncertainty, thereby including the often neglected procedural uncertainty, and derive theoretical expressions for them by usage of the classical bias-variance decomposition (Section 2). Subsequently, we discuss how epistemic uncertainty can be evaluated in a correct manner using statistical simulations, and we show with an illustrative example that the underlying model bias can lead to untrustworthy estimates if not taken into account, even if parts of the epistemic uncertainty are correctly estimated (Section 3). Furthermore, by assessing bias, data- and procedural uncertainty conceptually and in synthetic and real data experiments, we illustrate that common second-order uncertainty quantification methods only capture parts of the epistemic uncertainty, often confusing it with aleatoric uncertainty (Section 4).
2 Towards a fine-grained definition of epistemic uncertainty
For the scope of this paper, we define an instance space $\mathcal{X}\subset\mathbb{R}^{D}$, and a label or target space $\mathcal{Y}$, where $\mathcal{Y}\subseteq\mathbb{R}$ for the (univariate) regression, and $\mathcal{Y}=\{1,\dots,K\}$ for the classification case. We further assume having a dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ of size $N\in\mathbb{N}$, where each $(\boldsymbol{x}_{i},y_{i})$ is an independent and identically distributed (i.i.d.) realization of an unknown probability distribution $P$ on $\mathcal{X}\times\mathcal{Y}$. Let $p(y|\boldsymbol{x})$ be the associated (ground truth) conditional probability to observe label $y$ as an outcome given $\boldsymbol{x}$.
Assuming a parameter space $\Theta$ of the underlying probability distribution, for a given $\boldsymbol{x}\in\mathcal{X}$, we define $\boldsymbol{\theta}(\boldsymbol{x})\in\Theta$ as the parameters of $p(y|\boldsymbol{x})$.
One common representation of uncertainty, which is going to be the focus of this paper, is via a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ over these first-order parameters. Measures such as variance or entropy can then be used to estimate aleatoric and epistemic uncertainty. An estimate $\hat{p}(y|\boldsymbol{x})$ of the predictive distribution can then be obtained by Bayesian model averaging as
$\hat{p}(y|\boldsymbol{x})=\int_{\Theta}p(y|\boldsymbol{\theta},\boldsymbol{x})q(\boldsymbol{\theta}|\boldsymbol{x})\,d\boldsymbol{\theta}.$ (1)
2.1 The difficulty of defining epistemic uncertainty
Unlike aleatoric uncertainty, epistemic uncertainty does not have a universally accepted ground-truth value. Instead, it is inherently context dependent and must always be viewed relative to some benchmark. On the other hand, for different representations one can derive certain logical criteria that the aleatoric and epistemic uncertainty estimates have to fulfill. Empirically, Bengs et al. [27] argued that a second-order distribution should ideally be monotonic with respect to sample size and converge to a Dirac delta distribution if the dataset size goes to infinity. However, we argue that especially these empirical criteria are far from being fully established yet, and miss certain sources of epistemic uncertainty in practical applications, like procedural uncertainty [15], i.e. uncertainty due to randomness in the learning process. To this end, we provide a thorough taxonomy of the different sources of epistemic uncertainty.
As aleatoric uncertainty corresponds to the intrinsic randomness of the data generating process, one can directly associate it with the stochasticity induced by $p(y|\boldsymbol{x})$, using measures such as variance or entropy. Epistemic uncertainty, on the other hand, measures ”everything else” that is in principle reducible by collecting more knowledge, hence all uncertainty with respect to the parameters $\boldsymbol{\theta}$, but also the chosen hypothesis space $\mathcal{H}$. This uncertainty can stem from different sources, and several authors have discussed some of these sources [2, 15, 14]. Unifying those perspectives, we will distinguish in total four different sources of epistemic uncertainty (see Figure 1(b) for an overview). Having a loss function $\ell:\mathbb{P}(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ with associated expected loss $\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(\cdot,y)\big]$, a first important subdivision into three components can be made as follows:
  • Model uncertainty: arises when the hypothesis space $\mathcal{H}$ does not contain the true underlying distribution, i.e. $p(y|\boldsymbol{x})\mathrel{\vtop{\halign{#\cr\hfil$\in$\hfil\cr\hfil$/$\hfil\cr}}}\mathcal{H}$. In Figure 1(b) the notation $\hat{p}(y|\boldsymbol{x})$ is used to denote the Bayes-optimal model for a given hypothesis space. It can be formally defined as follows:
    $\hat{p}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{y|\boldsymbol{x}}\ell(\hat{p}(y|\boldsymbol{x}),y).$
  • Estimation uncertainty: quantifies how well the learning algorithm can approximate $\hat{p}(y|\boldsymbol{x})$, which is the best possible model given the hypothesis space.
  • Distributional uncertainty: is due to a shift in distribution between training and test data [14], respectively, and their associated conditional distributions $p(y|\boldsymbol{x})$ and $\tilde{p}(y|\boldsymbol{x})$.
Following recent work of Huang et al. [15], we further divide estimation uncertainty into data uncertainty and procedural uncertainty. To this end, we define $\mathcal{D}_{N}\sim P^{N}$ as the random variable reflecting the data uncertainty, where $P^{N}$ is the distribution over all datasets of size $N$. Similarly, we define $\gamma\sim P^{\gamma}$ as the random variable reflecting all procedural parameters (e.g. random weight initialization in Deep Ensembles), where $P^{\gamma}$ denotes the distribution over which these parameters are sampled. For a fixed dataset $\mathcal{D}_{N}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$, data uncertainty can then be formally defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and
$\hat{p}_{N}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\gamma}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}),y_{i})\big]\,$,
where $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ denotes a model trained on a given dataset of size $N$ and procedural parameters $\gamma$. Similarly, for a fixed set of procedural parameters, procedural uncertainty can be formally defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and
$\hat{p}_{\gamma}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{\hat{p}_{N,\gamma}(y|\boldsymbol{x})\in\mathcal{H}}\mathbb{E}_{\mathcal{D}_{N}}\big[\frac{1}{N}\sum_{i=1}^{N}\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}_{i}),y_{i})\big].$
Using this notation, estimation uncertainty is then defined as the difference between $\hat{p}(y|\boldsymbol{x})$ and $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$, as visualized in Figure 1(b).
2.2 Uncertainty from a generalized bias-variance perspective
From a statistician’s viewpoint, the classical bias-variance decomposition provides further insight into the interplay between estimation uncertainty and aleatoric uncertainty [20, 13]. While it is originally defined for the mean squared error as a loss function, it can also be generalized to any Bregman divergence [28, 29]. With $\ell:\mathbb{P}(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ being a proper scoring rule, for a given estimate $\hat{p}(y|\boldsymbol{x})$ of the conditional distribution, one can decompose the point-wise risk as
$\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(\hat{p}(y|\boldsymbol{x}),y)\big]$ $=\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(p(y|\boldsymbol{x}),y)\big]+\mathbb{E}_{y|\boldsymbol{x}}\big[\ell(p(y|\boldsymbol{x}),y)-\ell(\hat{p}(y|\boldsymbol{x}),y)\big]$
$=\underbrace{\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]}_{\text{aleatoric uncertainty}}+\underbrace{D_{\ell}\big(p(y|\boldsymbol{x}),\hat{p}(y|\boldsymbol{x})\big)}_{\text{epistemic uncertainty}}$ (2)
where $\mathbb{H}_{\ell}$ is the generalized Bayes risk of $\ell$ and $D_{\ell}$ is its associated divergence. For the empirical estimate, let $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ be the predictive distribution obtained with data uncertainty $D_{N}\sim P^{N}$ and procedural uncertainty $\gamma\sim P^{\gamma}$. Then the pointwise expected loss becomes [30]
$\mathbb{E}_{y|\boldsymbol{x}}\big[\mathbb{E}_{\mathcal{D}_{N},\gamma}[\ell(\hat{p}_{N,\gamma}(y|\boldsymbol{x}),y)]\big]=$ $\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]+\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(p(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))\big]$
$=$ $\underbrace{\mathbb{H}_{\ell}\big[p(y|\boldsymbol{x})\big]}_{\text{aleatoric unc.}}+\underbrace{D_{\ell}\big(p(y|\boldsymbol{x}),\bar{p}^{*}(y|\boldsymbol{x})\big)}_{\text{generalized bias}}$
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x})\big]}_{\text{generalized variance}},$ (3)
where $\bar{p}^{*}(y|\boldsymbol{x})=\operatorname*{\arg\min}_{p\in\mathbb{P}(\mathcal{Y})}D_{\ell}(p,\hat{p}_{N,\gamma}(y|\boldsymbol{x}))$ the dual mean of $\hat{p}_{N,\gamma}(y|\boldsymbol{x})$ under data- and procedural uncertainty. The first term in Eqn. 2.2 corresponds to the intrinsic variability within the data-generating process, and hence to aleatoric uncertainty. The last term reflects the generalized variance of the predictive distribution under data and procedural uncertainty, and ca,n under suitable conditions of the loss function $\ell$ [31] be further decomposed by means of the conditional bias-variance decomposition for Bregman divergences [30] as
$\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x})\big]$ $=\underbrace{\mathbb{E}_{\mathcal{D}_{N}}\big[D_{\ell}(\bar{p}^{*}(y|\boldsymbol{x}),\bar{p}_{|\mathcal{D}_{N}}^{*}(y|\boldsymbol{x}))\big]}_{\text{procedural uncertainty}}$
$+\underbrace{\mathbb{E}_{\mathcal{D}_{N},\gamma}\big[D_{\ell}(\bar{p}_{|\mathcal{D}_{N}}^{*}(y|\boldsymbol{x}),\hat{p}_{N,\gamma}(y|\boldsymbol{x}))\big]}_{\text{data uncertainty}}.$ (4)
Finally, the second term in Eqn. 2.2 reflects the generalised bias, which is the difference between the mean predicted distribution $\bar{p}^{*}(y|\boldsymbol{x})$ under data and model uncertainty and the true conditional distribution $p(y|\boldsymbol{x})$. For universal approximators, such as many types of deep neural networks, the model uncertainty is zero [32, 33, 34]. Such models are asymptotically unbiased, but in the finite data regime, a bias is unavoidable to make learning possible [20]. For more restrictive hypothesis spaces, the bias even does not disappear when the training dataset becomes infinitely large. Thus, in both cases, the bias term cannot be ignored when analysing epistemic uncertainty. We see that any proper scoring rule whose associated divergence can be related to a Bregman divergence allows for the split of its expected risk into an aleatoric and an epistemic part, where the epistemic part can be further split into a part representing bias, procedural and data uncertainty, respectively. In the next sections we will look at this decomposition more closely for the regression case, using mean squared error and negative log likelihood as loss functions.
3 Evaluating epistemic uncertainty via statistical simulations
3.1 The reference distribution
Quantifying epistemic uncertainty objectively is challenging, and evaluating the correctness of such quantifications is equally difficult. While calibration tests exist for aleatoric uncertainty in probabilistic models [35], epistemic uncertainty lacks established evaluation protocols and is often assessed indirectly through downstream tasks like out-of-distribution detection [5, 22, 36], robustness to adversarial attacks [37, 38], and active learning [39]. Correctness of second-order epistemic uncertainty representations can only be evaluated through statistical simulations, in which bias and variance can be analysed individually. To this end, Jürgens et al. [24] recently introduced the notion of a reference distribution, which is the theoretical distribution over first-order models if training data could be resampled. Here we further extend this notion by taking the variance due to procedural uncertainty into account:
Definition 1.
Consider a hypothesis space of models with $\phi\in\Phi$ model parameters and $\gamma\in\Gamma$ procedural parameters. We define the reference distribution w.r.t. $P^{N}$ and $P^{\gamma}$ as
$q_{N,\gamma}(\boldsymbol{\theta}|\boldsymbol{x},\phi):=\mathbb{P}(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})=\int_{\Gamma}\int_{\mathcal{X}\times\mathcal{Y}}\mathbb{I}\Big[(\boldsymbol{\hat{\theta}}_{N,\gamma}(x)=\boldsymbol{\theta})\Big]dP^{N}dP^{\gamma},$ (5)
where $\boldsymbol{\hat{\theta}}_{N,\gamma}$ is the empirical estimate of $\boldsymbol{\theta}$, obtained via empirical risk minimization.
Intuitively, the reference distribution describes the uncertainty in the first-order parameters, induced by having a limited dataset and procedural uncertainty due to different weight initializations. In simulation studies, one can estimate the reference distribution by drawing $n_{d}$ training datasets $\mathcal{D}_{N_{1}},\dots,\mathcal{D}_{N_{n_{d}}}$ and $n_{\gamma}$ procedural parameters $\gamma_{1},\dots,\gamma_{n_{\gamma}}$ from $P^{N}$ and $P^{\gamma}$, respectively, and calculating the empirical distribution function. The (generalized) variance of the reference distribution describes the variability due to limited data expressivity as well as the procedural variance, and can be decomposed by means of the decomposition in Eqn. 2.2. In addition, in statistical simulations, where the data-generating process $p(y|\boldsymbol{x})$ is known, the bias of the reference distribution can also be measured.
3.2 Bias-variance decomposition in the regression case
We start this section by deriving expressions for aleatoric and epistemic uncertainty estimates according to Section 2.2 for the regression case, which is the setting we focus on for the remainder of the paper. Regression has the advantage over classification that our key findings can be more easily and intuitively visualized, and has also not been studied as extensively as the classification case for uncertainty disentanglement.
We consider the following data-generating process: $y_{i}=f(\boldsymbol{x}_{i})+\epsilon_{i}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(\boldsymbol{x}_{i}))$ is data-dependent (heteroscedastic) noise, and $f$ is a deterministic function, such that $p(\cdot|\boldsymbol{x})$ forms the density of a Gaussian distribution with mean $f(\boldsymbol{x})$ and variance $\sigma^{2}(\boldsymbol{x})$. For the case where the loss function is the mean squared error loss, $\ell_{\text{MSE}}(\hat{y},y)=(\hat{y}-y)^{2}$, Eqn. 2.2 then becomes the classical bias-variance decomposition:
$\mathbb{E}_{y|\boldsymbol{x}}\big[\mathbb{E}_{\mathcal{D}_{N},\gamma}(\hat{y}-y)^{2}|\boldsymbol{x}\big]=\sigma^{2}(\boldsymbol{x})+\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})+\text{bias}^{2}\big[\hat{y}|\boldsymbol{x}\big],$ (6)
where the variance of the prediction $\hat{y}$ can be further split into procedural and data uncertainty by means of the law of total variance:
$\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})=\mathbb{E}_{\mathcal{D}_{N}}\big[\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big]+\text{Var}_{\mathcal{D}_{N}}(\mathbb{E}_{\gamma}[\hat{y}|\boldsymbol{x}]|\mathcal{D}_{N}).$ (7)
3.3 Simulating data and procedural uncertainty
Figure 2: Estimated procedural and data uncertainty of the reference distribution for increasing sample sizes. The reference distribution is estimated by resampling training data and procedural parameters, using $n_{\gamma}=10$ and $n_{d}=20$. Procedural and data uncertainty are calculated using the decomposition formula in Eqn. 2.2. Even when the model predictions (complete line) have an important deviation (bias) from the ground-truth expected value (dashed line), the epistemic uncertainty estimate decreases with more data.
In the following we illustrate how one can use the reference distribution in Def. 1 to calculate expressions for data- and procedural uncertainty in the regression case. To this end, we adapt the synthetic regression setting from Rossellini et al. [40]. Precisely, we generate datasets $\mathcal{D}_{N}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where
$x_{i}\sim Beta(1.2,0.5),\quad y_{i}=\sin(\frac{1}{5(x_{i}+0.16)^{3}})+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}(x_{i})),$ (8)
with $\sigma^{2}(x)=x^{4}$. The setup can be seen in Figure 1(a). This dataset highlights the differences between aleatoric and epistemic uncertainty: for low $x$ values, the complexity of the underlying data-generating function is high, while data is limited, leading to high epistemic uncertainty. On the other hand, for $x$ close to $1$, aleatoric uncertainty is high, while simultaneously having more data points and a less complex underlying function, implying low epistemic uncertainty. In order to estimate the reference distribution we train models with $n_{d}$ training datasets and $n_{\gamma}$ random weight initializations. As the underlying model class for obtaining the individual risk minimizers $\hat{p}_{\gamma,N}$, we use an MLP. Details about the experimental setup are discussed in Appendix A.1111The implementation for all our examples can be found in this GitHub repository.. Individual estimates for data- and procedural uncertainty can be obtained by the decomposition formula in Eqn. 2.2. The bias can be calculated as the absolute distance between the mean prediction and the underlying ground truth conditional expectation.
Figure 2 shows the obtained procedural and data uncertainty of the reference distribution for different training dataset sizes. First, we see that the total variance of the reference distribution decreases when sample size increases, supporting the claim that the reference distribution properly represents data uncertainty. Second, we see that the estimate of the procedural uncertainty also decreases with an increasing sample size, which is not unexpected, because initial procedural configurations become less important when more data is observed. Third, in the region with high epistemic uncertainty, the reference distribution has a very large bias, due to the underlying complexity of the data generating function. The bias is in fact so high that aleatoric uncertainty cannot be correctly estimated anymore. This last finding confirms our central position: Any claim to properly represent epistemic uncertainty must examine not only how different sources of uncertainty are modeled, but also whether bias is systematically distorting the partitioning of uncertainty. When the bias is high, estimates of both epistemic and aleatoric uncertainty can be completely misleading.
4 Which part of epistemic uncertainty do current methods capture?
Throughout this section we evaluate methods for the heteroscedastic regression setting using the classical variance-based split of the predictive variance as in Eqn. 7. We hereby focus on the mean squared error case of the more general pointwise decomposition in Section 2.2, which is the most commonly used decomposition for regression in practice. Departing from a Bayesian perspective, we first describe in this section how various second-order methods claim to obtain disentangled estimates for aleatoric and epistemic uncertainty. Secondly, we analyse two representative methods that approximate the posterior distribution over the parameters in more detail, illustrating that they tackle (and neglect) different sources of epistemic uncertainty. In the last part of this section as well as in Appendix A.1, we present experimental results for eight selected methods, highlighting that all of them have difficulties to model the different sources of epistemic uncertainty in a complete manner.
4.1 Second-order uncertainty quantification methods
Bayesian methods form a principled approach to quantify and interpret epistemic uncertainty by modelling a prior and posterior distribution over the model parameters and computing the second-order distribution via Eqn. 1. For the regression setup that we consider in this section, a common approach is to use Heteroscedastic Gaussian Processes (GP, [19, 41]), which place priors on the parameters $f(\boldsymbol{x})$ and $\sigma^{2}(\boldsymbol{x})$, thereby analytically computing the predictive distribution using two Gaussian processes. For Bayesian neural networks, which have gained popularity in recent years, the posterior and second-order distribution cannot be analytically computed, due to the intractability of $p(\boldsymbol{\phi}|\mathcal{D}_{N})\propto p(\mathcal{D}_{N}|\boldsymbol{\phi})p(\boldsymbol{\phi})$ induced by the hypothesis space [1]. Due to the impossibility of performing exact inference in Bayesian neural networks, one typically obtains $d$ different samples $\{\boldsymbol{\phi}_{j}\}_{j=1}^{d}$ from $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ by use of approximation techniques that aim to find a posterior distribution via sampling methods. Here, the second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ of Eqn. 1 is then approximated by Monte Carlo integration:
$q(\boldsymbol{\theta}|\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\phi}\sim p(\boldsymbol{\phi}|\boldsymbol{x})}\big[q(\boldsymbol{\theta}\mid\boldsymbol{x},\boldsymbol{\phi})\big]\approx\frac{1}{d}\sum_{j=1}^{d}q(\boldsymbol{\theta}\mid\boldsymbol{x},\boldsymbol{\phi}_{j}).$ (9)
In the experiments below, we employ three of the most common methods to obtain $\{\boldsymbol{\phi}_{i}\}_{j=1}^{M}$. Variational inference, (VI, [42, 43]) considers a variational (simpler) distribution $q(\boldsymbol{\phi})$ and optimizes the lower bound of the marginal likelihood (ELBO). Hamiltonian Monte Carlo (HMC) uses Hamiltonian dynamics to sample $\{\boldsymbol{\phi}_{j}\}$ from the true posterior. Laplace approximation (LA, [44]) builds a Taylor expansion around the maximum-a-posteriori estimate $\boldsymbol{\phi}_{\text{MAP}}$, yielding a Gaussian approximation. Ensemble methods are also often interpreted as performing approximate Bayesian inference by finding multiple modes of $p(\boldsymbol{\phi}|\mathcal{D}_{N})$ after training multiple independent models. [45, 46]. In this work, we consider three types of ensembles. Deep Ensembles (DE, [9]), train $d$ networks on the same dataset $\mathcal{D}_{N}$, but with different random initialization of the weights $\boldsymbol{\phi}$. Similarly, in Bootstrap Ensembles (BE), $M$ networks are trained using different weight initializations, but each network uses a bootstrap sample $\mathcal{D}^{(j)}\subseteq\mathcal{D}_{N}$. While Bootstrap Ensembles have been shown to perform worse than training $d$ networks using Deep Ensembles [9], we include them in our analysis to estimate the effect of data uncertainty on the epistemic uncertainty estimates. Monte Carlo Dropout (MCD, [47]) can also be understood as a Bayesian approximation, where one trains a single neural network with dropout rate $\rho$ and performs multiple stochastic forward passes at inference time with the same dropout rate. In a different direction, evidential deep learning methods learn directly the second-order distribution over the predictive parameters $\theta$. Deep evidential regression (DER, [12]) learns hyperparameters of a conjugate prior (e.g. Normal-Inverse-Gamma for the univariate normal distribution) directly using a single network trained by second-order loss minimization.
4.2 Uncertainty disentanglement
From a second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$, one can decompose its representation of uncertainty into an estimate of aleatoric and epistemic uncertainty. Proposed measures focus on entropy [3], distance to reference sets [10], or proper scoring rules [48, 49]. The most commonly used one for regression is arguably the variance-based decomposition [3, 10], which considers the variance of the prediction $\hat{y}$, obtained via model averaging as in Eqn. 1. The latter can be decomposed into (estimates of) an aleatoric and an epistemic part:
$\text{Var}(\hat{y}|\boldsymbol{x})=\underbrace{\mathbb{E}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta}\mid\boldsymbol{x})}\big[\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big]}_{\text{ aleatoric estimate}}+\underbrace{\text{Var}_{\boldsymbol{\theta}\sim q(\boldsymbol{\theta}\mid\boldsymbol{x})}\big(\mathbb{E}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big)}_{\text{epistemic estimate}}$ (10)
Here, the first term estimating aleatoric uncertainty describes the average predicted variance $\sigma^{2}$. The second term captures the variance of the mean prediction $f$ and hence an estimate of epistemic uncertainty. Note here that, unlike in the bias-variance decomposition in Eqn. 2.2, the bias in the prediction is not represented by this decomposition. Since any second-order distribution does not incorporate ground-truth knowledge, any systematic bias goes unaccounted. Moreover, second-order uncertainty quantification methods lack informativeness about which source of epistemic uncertainty they capture. In the next paragraph we illustrate these two concerns for two methods that serve as representative examples.
4.3 Two representative cases
Example 1 (Deep Ensembles purely capture procedural uncertainty.).
Let $\gamma\sim P^{\gamma}$ denote the random seed (weight initialization, data ordering etc.) that differentiates each ensemble member’s training, which in turn are trained on the same dataset $\mathcal{D}_{N}$. For a Deep Ensemble of size $d$, each ensemble member’s predictive parameters are indexed over the set of random seeds $\{\gamma_{1},\dots,\gamma_{d}\}$, obtaining a set of predictions $\{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})=(f_{\gamma_{i}}(\boldsymbol{x}),\sigma_{\gamma_{i}}^{2}(\boldsymbol{x}))\}_{i=1}^{d}$. The empirical second-order distribution $q(\boldsymbol{\theta}|\boldsymbol{x})$ can be approximated as a mixture of equally weighted Dirac measures
$q(\boldsymbol{\theta}|\boldsymbol{x})=\frac{1}{d}\sum_{i=1}^{d}\delta_{\boldsymbol{\theta}_{\gamma_{i}}(\boldsymbol{x})}(\boldsymbol{\theta}),$
[…]
$\text{Var}(\hat{y}|\boldsymbol{x})$ $=\frac{1}{d}\sum_{i=1}^{d}\sigma_{\gamma_{i}}^{2}(\boldsymbol{x})+\frac{1}{d}\sum_{i=1}^{d}(f_{\gamma_{i}}(\boldsymbol{x})-\hat{f}(\boldsymbol{x}))^{2}$
$\longrightarrow$ $\,\mathbb{E}_{\gamma}\big[\text{Var}(\hat{y}|\boldsymbol{x},\gamma)\big]+\text{Var}_{\gamma}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big),\,d\rightarrow\infty.$
Here, the first part corresponds to the aleatoric uncertainty estimate and the epistemic uncertainty estimate is given by
$\text{Var}_{\gamma}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]\big)=\mathbb{E}_{\gamma}\big[(\mathbb{E}_{x|y}[\hat{y}|\boldsymbol{x},\gamma]-\mathbb{E}_{\gamma}[\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\gamma]])^{2}\big]=\mathbb{E}_{\gamma}\big[(\hat{y}_{N,\gamma}(\boldsymbol{x})-\hat{y}_{N}(\boldsymbol{x}))^{2}\big],$
where $\hat{y}_{N,\gamma}(\boldsymbol{x}):=f_{\gamma}(\boldsymbol{x})$ corresponds to the prediction of an individual ensemble member trained with random initialization $\gamma$ on $\mathcal{D}_{N}$ and $\hat{y}_{N}(\boldsymbol{x}):=\hat{f}(\boldsymbol{x})$ of the aggregated ensembles. Hence, we see that the predicted epistemic uncertainty corresponds to the procedural uncertainty as illustrated in Figure 1(b). Model uncertainty through bias as well as data uncertainty are not represented.
While Deep Ensembles might explore multiple (local) optima of the training loss by using different random seeds [45], Bayesian approaches attempt to explicitly approximate the posterior over parameters, which can yield more principled estimates of epistemic uncertainty. However, the practical methods used to approximate the posterior distribution also face restrictions: Approximate inference methods like the Laplace approximation or variational inference often capture only a single local mode [45], employ simplifying assumptions [25], or can be computationally costly in high-dimensional weight spaces. Other issues arise with methods that try to estimate the second-order distribution directly via loss minimisation, such as Deep Evidential Regression.
Example 2 (DER may fail to provide faithful epistemic uncertainty estimates.).
Deep Evidential Regression [12] learns to predict the parameters of a Normal–Inverse-Gamma second–order distribution
$q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})=\mathrm{NIG}\!\bigl(\gamma(\boldsymbol{x}),\nu(\boldsymbol{x}),\alpha(\boldsymbol{x}),\beta(\boldsymbol{x})\bigr),$
which forms the conjugate prior of the normal distribution. The parameters $(\gamma,\nu,\beta,\alpha)$ are learned by minimising a second-order loss, i.e. the loss function one derives when doing Bayesian inference with the NIG prior, together with a regularization term called the evidence, which controls the uncertainty for data points far from the training set. Using the variance-based decomposition, one obtains analytical expressions for the aleatoric and epistemic uncertainty estimates, namely
$\mathbb{E}_{\boldsymbol{\theta}\sim q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})}[\text{Var}(\hat{y}|\boldsymbol{x},\boldsymbol{\theta})\big]=\frac{\beta}{\alpha-1},\quad\text{Var}_{\boldsymbol{\theta}\sim q_{\text{NIG}}(\boldsymbol{\theta}|\boldsymbol{x})}\big(\mathbb{E}_{y|\boldsymbol{x}}[\hat{y}|\boldsymbol{x},\boldsymbol{\theta}]\big)=\frac{\beta(\alpha-1)}{\nu}.$
However, a line of work has shown that the method has severe drawbacks in the optimization [27, 50], as well as the estimation of aleatoric and epistemic uncertainty [51], and the representation of the second-order distribution [24], making it more of a heuristic than a principled method.
4.4 Synthetic data illustration
Figure 3: Deep Ensemble estimates, obtained with different sample sizes $N\in\{50,100,500\}$. The top row shows the training data points (blue crosses), the true conditional mean (dashed line), the model predictions (complete line), the $95\%$ confidence interval for aleatoric uncertainty (purple), as well as the true $95\%$ confidence interval of the aleatoric uncertainty (gray). The bottom row shows the estimate of epistemic uncertainty (green).
To illustrate the limitations of existing uncertainty quantification methods in producing reliable disentangled estimates, we showcase the behavior of the eight methods described in Section 4.1 for the problem setting described in Section 2. This section focuses primarily on the results obtained using Deep Ensembles, but we emphasize that the results for the other seven methods are also consistent with Deep Ensembles and coherent with our argumentation. Results for the remaining seven methods and detailed implementation specifications are included in Appendix A.1.
Figure 3 illustrates that epistemic uncertainty estimates decrease monotonically with sample size, consistent with the behavior observed for the reference distribution. This trend confirms that, as the dataset grows, models become increasingly confident, yet not necessarily more accurate, particularly in difficult-to-learn regions. In these settings, even with abundant data, the model can remain overconfident in systematically biased predictions. Such behavior reveals a key limitation of current epistemic uncertainty measures and disentanglement frameworks: they fail to account for systematic bias, leading to an underestimation of epistemic uncertainty. This pattern is not unique to Deep Ensembles, but is consistently observed across other methods (see Figures 814), indicating that this limitation is intrinsic to the way epistemic uncertainty is commonly represented rather than a method-specific artifact. Bias is approximated as specified in Eqn. 2.2: datasets are repeatedly resampled from the true data‐generating process, models are trained on each sample, and the mean deviation of the predictions from the true conditional mean is computed. Comprehensive results for every method are presented in Table 1; the table shows that bias is frequently larger than –and largely uncorrelated with– the corresponding epistemic uncertainty estimates. These findings underscore our central position: any attempt to disentangle aleatoric from epistemic uncertainty must assess not only how each component is modeled, but also the extent to which systematic bias distorts their partitioning.
Figures 6(c) and 7 further illustrate the relationship between the epistemic uncertainty estimates and the procedural and data uncertainty components obtained from the reference distribution. When comparing these components, the relation shifts upward once data uncertainty is included alongside procedural uncertainty, indicating that most methods primarily capture the procedural component of epistemic uncertainty. This observation highlights the necessity of adopting a more fine-grained taxonomy to identify which parts of epistemic uncertainty are being represented, as well as the importance of the reference distribution in empirically validating these estimates. On the other hand, the aleatoric uncertainty estimates on the right‐hand side of the data support improve as the sample size increases, the behaviour on the left‐hand side remains problematic. Specifically, the model persistently overestimates aleatoric uncertainty. A plausible explanation is that the approximate symmetry of the observations around the learned conditional mean induces the model to misinterpret systematic deviation as inherent data noise. Consequently, this misestimation is not corrected with additional data; instead, the model becomes increasingly confident in its erroneous predictions and continues to interpret the discrepancy from the ground-truth conditional mean as aleatoric uncertainty. This behavior is illustrated in Figure 5(a), which compares the average estimated aleatoric uncertainty with the true aleatoric noise in regions of low and high model bias. The figure shows that in areas of high bias, the estimated aleatoric uncertainty diverges substantially from the true value, demonstrating that biased predictions systematically distort the disentanglement of aleatoric and epistemic uncertainty.
4.5 Real data experiments
In this section, we extend our analysis to the New York taxi trip duration dataset published in this Kaggle competition. We begin by describing how the reference distribution is estimated, followed by an explanation of how we approximate the ground-truth aleatoric uncertainty. Finally, we present the results obtained with the Deep Ensemble method, and refer to the appendix for the results of the rest of the methods. The dataset consists of taxi trip records from New York City, where the target variable corresponds to the trip duration and the feature space comprises 14 input variables. The training set contains $1,206,737$ observations, and the test set includes $230,406$ observations. Details about the dataset source and preprocessing steps are provided in Appendix A.2. This dataset is particularly suitable for evaluating uncertainty estimation methods for several reasons. First, its large sample size allows us to analyze how epistemic uncertainty behaves when the data-driven component becomes relatively small compared to the model and procedural components, allowing us to estimate data and procedural uncertainty via the reference distribution. In the same sense, a simulated estimate of ground truth aleatoric uncertainty can be obtained as we explain later. Second, since the task is formulated as a heteroscedastic regression problem, where trip durations exhibit high variability depending on factors such as traffic conditions or route choice, it provides a natural setting for disentangling aleatoric from epistemic uncertainty. Finally, the dataset’s scale and heterogeneity make it a realistic benchmark for assessing whether uncertainty quantification methods can remain well-calibrated and informative in complex, high-dimensional environments. As in the synthetic data experiments, we assume that the target variable follows a normal distribution, and the model outputs both the mean and variance of this distribution.
Reference distribution. To estimate the reference distribution as described in Section A.1, we partition the training dataset into $100$ disjoint subsets of $12,067$ observations each (i.e., $n_{d}=100$). For each subset, we train $n_{\gamma}=5$ multilayer perceptrons (MLPs) with different random weight initializations, resulting in a total of $500$ independently trained models. Each model produces predictive means $\hat{y}_{N,\gamma}(\boldsymbol{x})$, and by computing the empirical variance across these 500 predictions, we obtain an estimate of $\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})$ as defined in Eqn. 7. This procedure operationalizes the reference distribution introduced in Section 3: the pseudo-resampling across datasets ($\mathcal{D}_{N}$) captures data uncertainty while the random initializations account for procedural uncertainty. Given the large sample size of the dataset, the partitioning approach provides a computationally feasible approximation of repeated sampling from the underlying data-generating process, effectively emulating access to its stochastic variability. This design allows us to decompose the epistemic uncertainty component into its data and procedural parts in a controlled manner, while being in a real world data setup. Figure 4(c) shows the resulting decomposition of total epistemic variance into procedural and data components. We observe that procedural uncertainty dominates across most of the input domain. This outcome reflects the high capacity and non-convex optimization landscape of neural networks: even with abundant data, different random initializations can converge to distinct local minima, leading to substantial procedural variation. By contrast, data uncertainty is smaller, indicating that resampling subsets from such a large dataset induces only limited diversity in the fitted models. It is important to note that the relative magnitudes of these two components depend on the experimental configuration. Increasing the subset size $n_{d}$ would typically reduce data uncertainty, while enlarging $n_{\gamma}$ would yield a more precise estimate of procedural uncertainty.
Aleatoric uncertainty. To approximate the ground-truth aleatoric uncertainty, we train an ensemble of $n_{\gamma}=5$ multilayer perceptrons (MLPs) on the full training dataset of $1,206,737$ observations. Since aleatoric uncertainty represents the irreducible noise inherent to the data-generating process, training on the complete dataset minimizes the influence of data-driven and procedural components, yielding the most reliable estimate of this term attainable in practice. We take the mean predicted variance across the ensemble members as the reference estimate of aleatoric uncertainty. To assess the effect of limited data and model bias on uncertainty disentanglement, we then compare this reference estimate with those obtained when each model is trained using only $1\%$ of the available data. This comparison highlights how insufficient data coverage and higher epistemic uncertainty can distort the estimated aleatoric component, revealing the extent to which current methods conflate epistemic and aleatoric sources of uncertainty.
Deep Ensemble estimates. As discussed in Example 1 and in the synthetic data experiments, Deep Ensembles predominantly capture the procedural component of epistemic uncertainty. This limitation arises because ensemble diversity originates solely from random initialization and stochastic training dynamics, while all models are trained on the same dataset. Consequently, the variance across ensemble members reflects the sensitivity of the optimization process rather than uncertainty about the data distribution itself. This behavior is observed in Figure 4(b), where the total variance of the reference distribution, represented on the $y$-axis as the sum of data and procedural uncertainty, is systematically higher than the epistemic uncertainty predicted by the Deep Ensemble. Quantitatively, the epistemic uncertainty estimate underestimates the reference distribution variance in approximately $80\%$ of the predictive samples. This systematic underestimation indicates that the ensemble fails to account for the data-driven component of epistemic uncertainty, which captures how model predictions would vary under different training datasets drawn from the same data-generating process. To further examine this relationship, Figure 4(a) decomposes the reference distribution into its procedural and data uncertainty components. When comparing only the procedural component of the reference distribution with the epistemic uncertainty predicted by the Deep Ensemble, the discrepancy disappears. This alignment confirms that Deep Ensembles primarily quantify procedural uncertainty, that is, the uncertainty arising from stochasticity in optimization and model initialization, while largely neglecting the variability induced by finite data sampling.
In practical terms, this result reinforces the theoretical argument presented in Section 4.3: the epistemic uncertainty estimated by Deep Ensembles does not reflect the full reducible uncertainty in the model but only its procedural subcomponent. Consequently, downstream tasks that rely on a correct epistemic uncertainty estimation, such as active learning, model selection, or out-of-distribution detection, may be misled by an incomplete uncertainty representation. In the same sense, Figure 5(b) shows that in the region where the model is more incorrect, the difference between the estimated aleatoric uncertainty and the simulated ground truth aleatoric uncertainty is larger than in the region where the model is more accurate, showing how a high bias can distort the disentanglement of aleatoric and epistemic uncertainty. The findings from this real-world regression task, therefore, underscore the importance of distinguishing between procedural and data uncertainty when evaluating or designing second-order uncertainty quantification methods.
Figure 4: (a): Procedural and data uncertainty estimates from the reference distribution of the taxi trip duration dataset. The procedural uncertainty component is larger than the data uncertainty component. (b): Comparison between the epistemic uncertainty estimated by the Deep Ensemble and the epistemic uncertainty obtained from the reference distribution. The Deep Ensemble underestimates the epistemic uncertainty $80\%$ of the time. (c): Relation between the epistemic uncertainty estimated by the Deep Ensemble and the procedural uncertainty coming from the reference distribution. On average the Deep Ensemble’s estimate captures the procedural uncertainty part of the total epistemic uncertainty.
5 Discussion
Epistemic uncertainty is generally not well-defined and still very much a heuristic concept for machine learning models, leading to interpretability issues. As a first attempt to formalise a more fine-grained notion, we present a taxonomy that unifies the leading definitions of epistemic uncertainty in the literature, and present a simulation-based evaluation protocol for it. Although our illustrative simulations involve heteroscedastic regression problems, our key findings are also applicable to homoscedastic regression and classification settings. Our simulations include one synthetic data experiment and one real data experiment. Whilst the synthetic data experiment allows for resampling, bias calculation, and comparison with the true aleatoric uncertainty, the taxi data experiment requires some modeling design decisions, which might introduce bias in the result. With these experiments, we show that epistemic uncertainty estimates underestimate the uncertainty by capturing only some part of one of the components described in our taxonomy, and ignoring the model bias. More precisely, the evaluated methods seem to capture the procedural component, while being away from the data uncertainty. This underestimation also induces an overestimation of aleatoric uncertainty in regions where the predictive bias is larger. An accurate estimation of the epistemic uncertainty requires deriving more task-specific and fine-grained evaluation protocols of the estimates in practical applications, and assessing the model bias when estimating it. It might be argued that the goal of an epistemic uncertainty estimate is not to quantify bias. However, precisely by not doing so, the disentanglement of uncertainties becomes problematic. Our experiments show how the aleatoric uncertainty estimates can become largely incorrect in regions where the model error increases, undermining the disentanglement of the two uncertainties.
Appendix A Experimental details
A.1 Synthetic data experiments
Reference distribution approximation
To obtain an empirical approximation of the reference distribution, we sample $20$ different datasets. For each dataset, we train $10$ different MLP models with different random weight initializations. Each MLP uses the whole dataset. The MLP models have $4$ hidden layers with $100$ neurons each. We use a learning rate of 0.009 and train the models for $500$ epochs. We repeated the process for different training dataset sizes: $50,100,500$ with batch sizes $32,32,64$ respectively. Seeds for the datasets are $\{7,42,2024,123,999,50,100,150,200,250,300,350,400,450,500,550,600,650,700,750\}$.
Experimental details by method
All methods are evaluated for different training dataset sizes: $\{50,100,500\}$ with batch sizes $\{32,32,64\}$ respectively. The experiments are repeated for $5$ different seeds: $7,42,123,999$ and $2024$. To conduct the experiments, a hyperparameter tuning over relevant parameters was performed.
Deep Ensemble. Trained with an MLP of $4$ hidden layers with $100$ neurons each, learning rate of 0.009. Training epochs: $50,250,500$. Every member of the deep ensemble uses the whole dataset.
Bootstrap Ensemble. Trained as the Deep Ensemble, but each ensemble member uses a boostrap sample with $60\%$ of the training data.
Monte Carlo Dropout. A single MLP of $4$ hidden layers with $100$ neurons each is trained. A dropout rate of $0.2$ is used at training and inference time. A learning rate of 0.002 is used.
Variational Inference. This model is trained based on Blundell et al. [52]. An MLP of $4$ hidden layers with $100$ neurons each converted to a Bayesian neural network is used. The model is trained with a learning rate of $0.005$, $200$ burn-in epochs, a beta value of $5$, $10$ Monte Carlo samples at training time, and $500$ at test. The model is evaluated for $250$ and $500$ training epochs.
Laplace Approximation. An MLP of $4$ hidden layers with $100$ neurons each is used to find the MAP estimate, together with a learning rate of $0.005$. For the Laplace step, a prior precision of $1$ and a noise value of $2$ are used. The model is trained for $200,400$, and $800$ epochs, and uses $1000$ samples from the posterior distribution. When using a training sample size of $500$, a batch size of 128 is used.
Hamiltonian Monte Carlo. An MLP of $4$ hidden layers with $100$ neurons each is used as the underlying model. The sampling step consists of $200$ samples with a sampling step of $0.00015$. The first $50$ samples are always burnt. A $\tau$ value of $1$ is used as the indicator for a bendy function. At inference time, we sample $1000$ times from the posterior distribution and drop the first $50$. The model is trained for $1000,2000$, and $3000$ epochs. When using a sample size of 500, a batch size of 128 is used.
Deep Evidential Regression. Leaning on the setup of Amini et al. [12], an MLP of $4$ hidden layers with $100$ neurons each is used as the underlying model. The model is trained with a learning rate of $0.0003$ over $5000$ epochs, and a regularization strength of $\lambda=0.01$.
Heteroscedastic Gaussian Process. Leaning on the variational-sparse recipe of Titsias [41], we use the GPytorch ApproximateGP module to train a two-output GP, trained with an RBF kernel. The number of inducing points (see [20] for an explanation) is set to $256$, the learning rate to 0.005, epochs to $2000$.
Further experimental results
In this section, we present additional experimental results for the other methods we tested. The data is generated according to the setting described in Section 2. We conduct the same experiment as described for the Deep Ensemble case in Section 4.4. The hyperparameters of each method were chosen according to Section A. The results can be seen in Figure 8-14. We highlight two additional key observations: First, we observe that Bayesian approaches –specifically Laplace Approximation (LA) and Hamiltonian Monte Carlo (HMC)– require a larger amount of data to produce accurate estimates of aleatoric uncertainty in the right-hand region of the input space. This behavior aligns with theoretical expectations of Bayesian inference: The increased data requirement reflects the greater difficulty these models encounter in learning the underlying data-generating function, which in turn leads to higher epistemic uncertainty–consistent with models that make inaccurate predictions. Second, the Deep Evidential Regression (DER) method consistently produces poor estimates of aleatoric uncertainty across all sample sizes and throughout the entire input domain. These inaccuracies are accompanied by overconfident predictions, as evidenced by the systematic underestimation of epistemic uncertainty. This flawed characterization of uncertainty is consistent with findings reported in recent literature [24]. We see that every method with a low epistemic uncertainty estimate overestimates the aleatoric uncertainty in the left-hand side of the figure due to their inability to acknowledge the bias as part of their epistemic uncertainty representations. Also, for more data-intensive methods, such as the Bayesian approaches, a more careful and realistic estimate of the epistemic uncertainty is obtained.
(a) Synthetic data (Equation 8)
(b) Taxi dataset.
Figure 5: Comparison of the estimated aleatoric uncertainty with the true (simulated) aleatoric noise across regions characterized by different levels of model bias and inherent data noise. The notation $\sigma_{i}$ for $i\in{1,2,3,4}$ denotes the true aleatoric uncertainty for each of the four regions: 1. low bias and low true noise, 2. high bias and low true noise, 3. low bias and high true noise, and 4. high bias and high true noise. For the synthetic dataset, the division between high- and low-noise regions follows Eqn. 8: points with $x<0.6$ are considered low-noise, while those with $x\geq 0.6$ are considered high-noise. Within each noise region, samples are ranked according to their model bias, and the lower half of points (by bias magnitude) defines the low-bias region, while the upper half defines the high-bias region. For the taxi dataset, the high- and low-noise regions are determined using the $90$-th percentile of the simulated true aleatoric uncertainty, as described in Section 4.5. Within each noise region, data points are further divided into low- and high-bias subsets based on the 70th percentile of the estimated model bias. This construction allows for a direct comparison between the estimated and true aleatoric uncertainty across systematically defined regions, highlighting how model bias affects the reliability of aleatoric uncertainty estimates in both synthetic and real-world scenarios.
Figure 6: Epistemic uncertainty estimates for the synthetic dataset. (a): Decomposition of the epistemic uncertainty estimated with the reference distribution into procedural and data uncertainties. There exist a strong correlation between both, which is expected, and procedural uncertainty overweighs across almost all the data points. (b): Comparison between the epistemic uncertainty estimated with the reference distribution (both data and procedural components) and the epistemic uncertainty estimated with the Deep Ensemble. Here, the epistemic part from the reference distribution is larger, showing that the Deep Ensemble underestimates this uncertainty. (c): When the data uncertainty component is removed from the estimate of the reference distribution, both estimates are much closer, suggesting that the Deep Ensemble fails to capture the data uncertainty component of the epistemic uncertainty.
(a) Bootstrap
(b) Bootstrap
(c) GP
(d) GP
(e) EDL
(f) EDL
(g) Laplace
(h) Laplace
(i) MC Dropout
(j) MC Dropout
(k) VI
(l) VI
Figure 7: Comparison of procedural and data uncertainties vs epistemic uncertainty across methods on the Sine dataset.
Figure 8: VI estimates, obtained with different sample sizes as in Figure 3.
Figure 9: GP estimates, obtained with different sample sizes as in Figure 3.
Figure 10: Bootstraped ensemble estimates, obtained with different sample sizes as in Figure 3.
Figure 11: MC Dropout estimates, obtained with different sample sizes as in Figure 3.
Figure 12: HMC estimates, obtained with different sample sizes as in Figure 3.
Figure 13: Laplace estimates, obtained with different sample sizes as in Figure 3.
Figure 14: DER estimates, obtained with different sample sizes as in Figure 3.
Training sample size: $N=100$
Model $\hat{\sigma}(x)$ $\text{Var}\big[\hat{y}|\boldsymbol{x}\big]$ $\mathbb{E}_{\mathcal{D}_{N}}[\hat{y}(x)-f(x_{i})]$ $\hat{\sigma}(x)-\sigma(x)$
Reference 0.83 (0.01) 0.041 (0.005) 0.084 (0.002) 0.248 (0.004)
Bootstrap 0.87 (0.167) 0.006 (0.008) 0.125 (0.022) 0.239 (0.027)
DeepEnsemble 0.87 (0.148) 0.007 (0.011) 0.147 (0.018) 0.235 (0.021)
EDL 0.21 (0.002) 0.003 (0.0) 0.269 (0.01) 0.751 (0.005)
ViELBONet 0.95 (0.064) 0.079 (0.001) 0.127 (0.021) 0.234 (0.022)
MCDropoutNet 0.90 (0.097) 0.009 (0.002) 0.190 (0.026) 0.334 (0.033)
HMCNet 9.08 (13.734) 2.131 (0.224) 0.179 (0.015) 7.870 (0.853)
LaplaceNet 2.43 (5.559) 4.211 (3.047) 0.133 (0.024) 0.800 (0.540)
GP 0.86 (0.048) 0.008 (0.001) 0.194 (0.013) 0.308 (0.020)
Training sample size: $N=500$
Model $\hat{\sigma}(x)$ $\text{Var}\big[\hat{y}|\boldsymbol{x}\big]$ $\mathbb{E}_{\mathcal{D}_{N}}[\hat{y}(x)-f(x_{i})]$ $\hat{\sigma}(x)-\sigma(x)$
Reference 1.04 (0.016) 0.013 (0.001) 0.044 (0.001) 0.076 (0.003)
Bootstrap 0.98 (0.100) 0.003 (0.004) 0.085 (0.001) 0.131 (0.010)
DeepEnsemble 0.99 (0.136) 0.004 (0.005) 0.100 (0.002) 0.122 (0.010)
EDL 0.67 (0.247) 0.013 (0.007) 0.168 (0.016) 0.967 (0.031)
ViELBONet 1.06 (0.115) 0.009 (0.000) 0.097 (0.005) 0.122 (0.012)
MCDropoutNet 1.06 (0.079) 0.008 (0.003) 0.119 (0.006) 0.163 (0.008)
HMCNet 4.17 (20.356) 0.378 (0.017) 0.129 (0.003) 1.936 (1.495)
LaplaceNet 1.29 (3.777) 0.487 (3.802) 0.115 (0.016) 0.187 (0.345)
GP 1.04 (0.069) 0.008 (0.003) 0.172 (0.005) 0.152 (0.014)
Table 1: Aleatoric and epistemic uncertainty estimates, bias, and distance from the aleatoric uncertainty to ground truth noise across models. Values are means over the test set from five runs; standard deviations are in parentheses.
In Table 1, we calculate the average aleatoric and epistemic uncertainty estimates for the eight different methods we analyse as showcases for our arguments. The last two columns show the average bias with respect to the ground truth conditional mean and the mean absolute distance from the aleatoric uncertainty estimate to the ground truth noise, respectively. It shows how methods that become overconfident with more training data are highly biased towards the true conditional mean. Furthermore, no correlation between epistemic uncertainty and bias is observed, showing how the unfaithfulness of the estimation is not only due to the order of magnitude of the estimands but also to the lack of information about the systematic deviations from the ground truth.
A.2 Real data experiments
Data preprocessing
The raw dataset used for this task can be found in this Kaggle competition. The goal is to predict the duration of a taxi trip in New York. The dataset contains $1,458,644$ observations from 2016. The code to preprocess the data is in the GitHub repository of this paper. The target variable is transformed according to $y^{\prime}=\log(y+1)$.
Data splitting. We consider the data before June as the training dataset and the data after June as the test dataset. In this way, we get $1,206,737$ training points and $230,406$ test points. However, only $1\%$ of the dataset is used for training the models. The simulated ground truth aleatoric uncertainty is obtained by using the complete dataset as explained in Section 4.
Models
Deep Ensemble. Trained with an MLP of $10$ hidden layers with $300$ neurons each, learning rate of $0.0009$. Training epochs: $50$. Every member of the deep ensemble uses the whole dataset (containing $1\%$ of the total training dataset).
Bootstrap Ensemble. Trained as the Deep Ensemble, but each ensemble member uses a boostrap sample with $60\%$ of the training data (out of the $1\%$).
Monte Carlo Dropout. A single MLP of $10$ hidden layers with $300$ neurons each is trained. A dropout rate of $0.2$ is used at training and inference time. A learning rate of $0.0009$ is used.
Variational Inference. This model is trained based on Blundell et al. [52]. An MLP of $10$ hidden layers with $300$ neurons each converted to a Bayesian neural network is used. The model is trained with a learning rate of $0.0009$, 2020 burn-in epochs, a beta value of $10$, $10$ Monte Carlo samples at training time, and $500$ at test. The model is evaluated for $50$ training epochs.
Laplace Approximation. An MLP of $4$ hidden layers with $100$ neurons each is used to find the MAP estimate, together with a learning rate of $0.0009$. For the Laplace step, a prior precision of $1$ and a noise value of $0.1$ are used. The model is trained for $50$ epochs, and uses $100$ samples from the posterior distribution.
Hamiltonian Monte Carlo. An MLP of $10$ hidden layers with $300$ neurons each is used as the underlying model. The sampling step consists of $70$ samples with a sampling step of $0.00015$. The first $10$ samples are always burnt. A $\tau$ value of $1$ is used. At inference time, we sample $100$ times from the posterior distribution and drop the first $20$. The model is trained for $50$ epochs.
Heteroscedastic Gaussian Process. The implementation is the same as in the synthetic data case. The number of inducing points (see [20] for an explanation) is set to $256$, the learning rate to $0.0009$, epochs to $50$.
A.2.1 Further experimental results
In Section 4.5 we discuss how the Deep Ensembles capture only the procedural part of the epistemic uncertainty, effectively underestimating it. In Figure 15 we can see how the Bootstrap and Monte Carlo Dropout methods have a very similar behavior: a strong correlation between the epistemic uncertainty and the procedural uncertainty, that is shifted when the data uncertainty component is considered.
Model $\hat{\sigma}^{2}$ $\hat{\sigma}^{2}(x)-\sigma_{*}^{2}(x)$ $\mathrm{Var}\!\big[\hat{y}|\boldsymbol{x}\big]$ $\mathrm{Var}_{\mathcal{D}_{N},\gamma}\!\big(\hat{y}|\boldsymbol{x}\big)$- $\mathrm{Var}\!\big[\hat{y}|\boldsymbol{x}\big]$ $\mathbb{E}_{\mathcal{D}_{N}}\!\big[\mathrm{Var}_{\gamma}\!\big(\hat{y}|\boldsymbol{x}\big)|\mathcal{D}_{N}\big]$- $\mathrm{Var}\!\big[\hat{y}|\boldsymbol{x}\big]$
DE 0.1222 0.0004 0.0083 0.0046 0.0013
Bootstrap 0.1127 0.0091 0.0082 0.0047 0.0014
MC Dropout 0.1284 0.0066 0.0106 0.0023 0.001
VI 1.8859 1.7641 0.0055 0.0074 0.0041
Laplace 0.1754 0.0536 0.0137 0.0008 0.0041
HMC 0.2319 0.1101 0.1148 0.1019 0.1052
GP 1.0334 0.9116 0.0636 0.0507 0.054
Table 2: Taxi dataset uncertainties. $\text{Var}\big[\hat{y}|\boldsymbol{x}\big]$ represents the epistemic uncertainty estimated by each model. $\text{Var}_{\mathcal{D}_{N},\gamma}(\hat{y}|\boldsymbol{x})$ corresponds to the epistemic uncertainty obtained with the reference distribution, while $\mathbb{E}_{\mathcal{D}_{N}}\big[\text{Var}_{\gamma}(\hat{y}|\boldsymbol{x})|\mathcal{D}_{N}\big]$ is the procedural uncertainty obtained by the reference distribution. The values are the average over all test data points.
(a) Bootstrap
(b) Bootstrap
(c) GP
(d) GP
(e) HMC
(f) HMC
(g) Laplace
(h) Laplace
(i) MC Dropout
(j) MC Dropout
(k) VI
(l) VI
Figure 15: Comparison of procedural and data uncertainties vs epistemic uncertainty across methods on the Taxi dataset. Figures in purple (first figure of every method) show the relation between the epistemic uncertainty from the reference distribution (including procedural and data uncertainties), and green pictures (second figure of every method) show the relationship between the procedural uncertainty from the reference distribution and the epistemic uncertainty from each method. In general, we see that adding the data uncertainty component shifts the relation between the two estimates.